### v3

# Maria Bazotte, Margarida Carvalho, and Thibaut Vidal

[…]
Maria Bazotte, Margarida Carvalho, and Thibaut Vidal
[…]
Solving Two-Stage Programs with Endogenous Uncertainty via Random Variable Transformation
[…]
Solving Two-Stage Stochastic Programs with Endogenous Uncertainty via Random Variable Transformation
[…]
Maria Bazotte
[…]
Margarida Carvalho
[…]
Thibaut Vidal
[…]
Real-world decision-making problems often involve decision-dependent uncertainty, where the probability distribution of the random vector depends on the model decisions. Few studies focus on two-stage stochastic programs with this type of endogenous uncertainty, and those that do lack general methodologies. We propose a general method for solving a class of these programs based on random variable transformation, a technique widely employed in probability and statistics. The random variable transformation converts a stochastic program with endogenous uncertainty (original program) into an equivalent stochastic program with decision-independent uncertainty (transformed program), for which solution procedures are well-studied. Additionally, endogenous uncertainty usually leads to nonlinear nonconvex programs, which are theoretically intractable. Nonetheless, we show that, for some classical endogenous distributions, the proposed method yields mixed-integer linear or convex programs with exogenous uncertainty. We validate this method by applying it to a network design and facility-protection problem, considering distinct decision-dependent distributions for the random variables. While the original formulation of this problem is nonlinear nonconvex for most endogenous distributions, the proposed method transforms it into mixed-integer linear programs with exogenous uncertainty. We solve these transformed programs with the sample average approximation method. We highlight the superior performance of our approach compared to solving the original program in the case a mixed-integer linear formulation of this program exists.
[…]
decision-dependent or endogenous uncertainty, network design and facility protection, random variable transformation, two-stage stochastic programs
1 Introduction
Uncertainty is an inherent characteristic of most real-world decision-making problems in fields such as network design planning (Contreras et al. 2011), vehicle routing (Gendreau et al. 1996), and scheduling (Restrepo et al. 2017). Stochastic programming is a well-known framework in mathematical programming to model optimization problems with uncertain parameters; see, e.g., (Birge and Louveaux 2011) and (Shapiro et al. 2021) for a summary of the theoretical aspects and the most popular algorithms. Research in this area has focused mainly on exogenous or decision-independent uncertainty (Li and Grossmann 2021), where the random vector is independent of the model decisions. In contrast, endogenous or decision-dependent uncertainty, where the random vector depends on the decisions, has received less attention, although it is prevalent in critical applications such as pre-disaster supply chain management (Peeta et al. 2010, Bhuiyan et al. 2020), production planning (Ekin 2018), and R&D portfolio selection (Solak et al. 2010).
In the context of endogenous uncertainty, different types of decision dependency may occur (Jonsbrten et al. 1998). Goel and Grossmann (2006) proposed classifying decision dependency into two types: For problems with Type 1 endogenous uncertainty or decision-dependent probability, decisions impact the underlying probability distribution of the random parameters. Although little research focuses on this type of uncertainty, as noted by Li and Grossmann (2021), it deals with various problems of practical relevance, such as the design of communication networks, power plants, and warehouses and resource allocation in queuing systems (Pflug 1990). With Type 2 endogenous uncertainty, also called “exogenous uncertainty with endogenous observability” (Mercier and Van Hentenryck 2011), decisions influence the time at which the realization of the random variables is totally or partially revealed. Thus, with Type 2 endogenous uncertainty, the scenario tree depends on the decisions (Jonsbrten et al. 1998). Stochastic programs with Type 2 uncertainty have received more attention than those with Type 1, being mostly multistage programs.  Apap and Grossmann (2017) and Li and Grossmann (2021) provide concise reviews of methodologies and algorithms for multistage programs with Type 2 endogenous uncertainty. In the present work, we focus on Type 1 endogenous uncertainty, in particular, on two-stage stochastic mixed-integer linear programs with decision-dependent probabilities (2-SMIPD), in which the probability distribution depends on the first-stage decisions. In what follows, we refer to Type 1 endogenous uncertainty as endogenous uncertainty.
The mathematical properties of a stochastic program, such as convexity, are strongly linked to whether its uncertainty is modeled as endogenous or exogenous. Consequently, solution procedures typically employed for stochastic programs with exogenous uncertainty become complex when dealing with endogenous uncertainty, necessitating a priori a comprehensive analysis of its underlying properties (Varaiya and Wets 1989). In fact, endogenous uncertainty usually leads to nonlinear nonconvex programs (Dupacová 2006), which are theoretically intractable (NP-hard). Moreover, few studies focus on general-purpose methodologies to solve programs with endogenous uncertainty sharing common properties (i.e., methodologies are typically problem-specific).
Contributions.
The main contribution of this work is a general method based on random variable transformation to address 2-SMIPD with discrete or continuous endogenous random variables. This method resolves two main issues in the literature: the lack of generalist methods and the difficulty of solving the usual nonconvex nonlinear programs with endogenous uncertainty. More specifically, the contributions of this work are as follows: (i) We define a transformation function to model the endogenous uncertainty by combining first-stage decisions with decision-independent random variables, thereby converting stochastic programs with endogenous uncertainty (i.e., original programs) into stochastic programs with exogenous uncertainty (i.e., transformed programs). (ii) We identify different techniques to define appropriate transformation functions to deal with endogenous random variables following classical decision-dependent probability distributions, including a commonly studied case known as the “discrete selection of distributions”. (iii) We show that, for these distributions, the obtained stochastic programs with exogenous uncertainty have a convex continuous relaxation. (iv) We investigate the implications of endogenous uncertainty when applying sampling methods, specifically the sample average approximation (SAA) method, to the transformed program. (v) We apply the proposed approach to the network design and facility protection problem (NDFPP), considering different decision-dependent distributions. (vi) Finally, for all endogenous distributions, we analyze the statistical estimators of the SAA applied to the transformed program of the NDFPP. In less than 1 hour on average, the SAA applied to the transformed program finds feasible solutions with significantly small over-estimated optimality gaps (between $-1\%$ and $1\%$) and a high confidence interval for all decision-dependent distributions. For the NDFPP with most endogenous distributions, the deterministic equivalent program (DEP) of the original program is nonlinear and nonconvex. The only exception is the discrete selection of distributions case, for which the DEP of the original program is a mixed-integer linear program (MILP) obtained by auxiliary linearizations (Bhuiyan et al. 2020). In a considerably shorter time, the SAA applied to the transformed program obtains a better over-estimator of the optimality gap than the optimality gap obtained by solving this mixed-integer DEP (a difference of around $24\%$ on average). Solving the transformed program with the SAA also obtains better solutions than when using an average scenario, i.e., the expected value problem (EV).
Organization.
The remainder of this work is structured as follows: Section 2 formalizes the class 2-SMIPD and discusses related works on endogenous uncertainty. Section 3 presents our general methodology to obtain the transformed program and discusses the application of methods such as the SAA to solve this resulting program. Section 4 catalogs techniques to define transformations to various endogenous distributions, and Section 5 explains how to apply this approach to the NDFPP. Finally, Section 6 reports the experimental results, comparing the proposed method with other solution procedures, and Section 7 concludes our analyses.
2 Problem Definition and Related Works
We focus on two-stage stochastic mixed-integer linear programs with decision-dependent probabilities (2-SMIPD). Let $x\in\mathbb{R}^{n_{1}}_{+}$ be the vector of first-stage variables (i.e., the “here-and-now” decisions) and $\xi(\omega)\in\mathbb{R}^{U}$ a random vector that depends on the random events $\omega\in\Omega$. For the sake of simplicity, we also denote $\boldsymbol{\xi}=\xi(\omega)$ and designate $\xi$ as a realization of this random vector. We use $\mathbb{E}_{\boldsymbol{\xi}\mid x}[\cdot]$ to denote the expected value operator with respect to $\boldsymbol{\xi}$ given the first-stage variables $x$. Finally, $y(\omega)\in\mathbb{R}^{n_{2}}_{+}$ is the vector of second-stage variables given a realization of the random events. The standard mathematical formulation for the 2-SMIPD is
$\min$ $c^{\top}x+\mathbb{E}_{\boldsymbol{\xi}\mid x}\left[Q\boldsymbol{(}x,\xi(\omega)\boldsymbol{)}\right]$ (1a)
s.t. $Ax\geq b$ (1b)
$x\in\mathcal{X}$ (1c)
where $\textrm{where }Q\boldsymbol{(}x,\xi(\omega)\boldsymbol{)}=\min\{q(\omega)^{\top}y(\omega):T(\omega)x+W(\omega)y(\omega)\geq h(\omega),y(\omega)\in\mathcal{Y}\}.$ (1d)
In Program (1), $\mathcal{X}\subseteq\mathbb{R}^{n_{1}}_{+}$ and $\mathcal{Y}\subseteq\mathbb{R}^{n_{2}}_{+}$ are two compact sets imposing non-negativity and continuous, integer, or binary restrictions. The first stage has the matrices $A$, $b$, and $c$ with appropriate dimensions. The second-stage Program (1d) is established after the realization of the random events $\omega\in\Omega$ and has the matrices $q(\omega)$, $h(\omega)$, $T(\omega)$, and $W(\omega)$ with possible random elements and appropriate dimensions. Thus, the random vector $\xi^{T}(\omega)$ combines these random elements, i.e., $\xi^{T}(\omega)=\boldsymbol{(}q(\omega)^{\top},h(\omega)^{\top},T(\omega),W(\omega)\boldsymbol{)}$, and its support (according to its unconditional distribution on $x$) is denoted by $\Xi\subseteq\mathbb{R}^{U}$. For the sake of clarity, we define the random vectors $\boldsymbol{\xi}_{x}^{\top}=\boldsymbol{(}q_{x}(\omega)^{\top},h_{x}(\omega)^{\top},T_{x}(\omega),W_{x}(\omega)\boldsymbol{)}$ for $x\in\mathcal{X}$, which follow the conditional probability distribution of the random vector $\boldsymbol{\xi}$ given $x$ and have support $\Xi_{x}\subseteq\Xi$. Finally, we set $g(x)=\mathbb{E}_{\boldsymbol{\xi}\mid x}[Q\boldsymbol{(}x,\xi(\omega)\boldsymbol{)}]$.
The 2-SMIPD assumes that the probability distribution of $\boldsymbol{\xi}$ depends on the first-stage decisions $x\in\mathcal{X}$ (i.e., it computes the conditional expectation $\mathbb{E}_{\boldsymbol{\xi}|x}[\cdot]$ instead of the standard expectation $\mathbb{E}_{\boldsymbol{\xi}}[\cdot]$ used in the exogenous case). This endogenous probability influences the properties of the resulting formulation. For example, unlike the exogenous case, the convexity of the continuous relaxation of $Q\boldsymbol{(}x,\xi(\omega)\boldsymbol{)}$ in Program (1) does not necessarily yield the convexity of the continuous relaxation of $\mathbb{E}_{\boldsymbol{\xi}|x}[Q\boldsymbol{(}x,\xi(\omega)\boldsymbol{)}]$ (Varaiya and Wets 1989, Dupacová 2006). As a result, algorithms commonly used for solving two-stage programs with exogenous uncertainty may not be suitable for 2-SMIPD.
This work assumes that the conditional probability distribution is known a priori and defined by a distinct probability mass or density function for $x\in\mathcal{X}$ (i.e., $x$ is a parameter of the distribution). We justify this assumption based on its broad application in the literature, as reviewed in Section 2.2. We consider it a reasonable extension of the usual stochastic optimization assumption for programs with exogenous uncertainty (i.e., random vectors with known probability distributions).
For the sake of simplicity, we next discuss our reasoning, assuming that the random vector $\boldsymbol{\xi}$ is discrete and has finite support $\Xi$; the process is analogous for the continuous case. We write the DEP in the extensive form of Program (1):
$\min$ $c^{\top}x+\sum_{k\in K}p_{k}(x)q_{k}^{\top}y_{k}$ (2a)
[…]
$T_{k}x+W_{k}y_{k}\geq h_{k}$ $\forall\ k\in K$ (2b)
$y_{k}\in\mathcal{Y}$ $\forall\ k\in K,$ (2c)
where $K$ is the set of realizations of $\boldsymbol{\xi}$ (i.e., the set of scenarios defined by the support $\Xi$). For each scenario $k\in K$, $q_{k}$, $h_{k}$, $T_{k}$, and $W_{k}$ have appropriate dimensions, and $\xi_{k}^{\top}=(q_{k}^{\top},h_{k}^{\top},T_{k},W_{k})$. In addition, $p_{k}(x):\mathcal{X}\rightarrow[0,1]$ is a function returning the conditional probability of scenario $k\in K$ given the first-stage variables $x$, in other words, $\boldsymbol{\xi}_{x}$ follows the probability distribution defined by $p_{k}(x)$ for every $k\in K$. We remark that scenarios in $\Xi\backslash\Xi_{x}$ have zero probability according to $p_{k}(x)$.
2.1 Background of Decision-Dependent Probability Distribution
In Program (2), the probability function $p_{k}(x)$ multiplies the second-stage objective term $q_{k}^{\top}y_{k}$, leading to nonlinear terms. Handling this nonlinearity requires describing $p_{k}(x)$, which depends on the assumptions of $\boldsymbol{\xi}$. When the $U$ elements of the random vector $\boldsymbol{\xi}$ are independent, we have
$p_{k}(x)=\prod_{u=1}^{U}\hat{p}_{\hat{r}(u,k)}^{u}(x)\ \forall\ k\in K,$ (3)
where $\hat{r}(u,k)$ returns the realization in $\mathcal{R}_{u}=\{1,\ldots,R_{u}\}$ of a random element $u\in\mathcal{U}=\{1,\ldots,U\}$ in scenario $k\in K$. In addition, $\hat{p}_{\hat{r}(u,k)}^{u}(x):\mathcal{X}\rightarrow[0,1]$ is a function on $x$ returning the conditional probability of the realization $\hat{r}(u,k)$ of the random element $u$ in scenario $k\in K$. The function $p_{k}(x)$ can assume other configurations that do not necessarily use $\hat{p}_{\hat{r}(u,k)}^{u}(x)$ but instead return the conditional probability of the whole random vector $\boldsymbol{\xi}$.
A commonly studied case where decisions consist of choosing a probability distribution among a predefined set of alternative distributions is designated the “discrete selection of distributions” (Dupacová 2006, Hellemo et al. 2018) and has been extensively studied. In this case, the decision-dependent probabilities are represented by a probability distribution for each independent random element $u\in\mathcal{U}$ among a set $\mathcal{D}_{u}=\{1,\ldots,D_{u}\}$ of available probability distributions, which could be discrete or continuous. For discrete endogenous random variables, $p_{k}(x)$ are given by
$p_{k}(x)$ $=\prod_{u\in\mathcal{U}}\left(\sum_{d\in\mathcal{D}_{u}}\overline{p}_{\hat{r}(u,k)}^{ud}x_{u}^{d}\right)\ \forall\ k\in K$ (4)
$\sum_{d\in\mathcal{D}_{u}}x_{u}^{d}$ $=1\ \forall\ u\in\mathcal{U},$ (5)
where $x_{u}^{d}\in\{0,1\}$ is a subvector of $x\in\mathcal{X}$ representing the chosen distribution. The parameter $\overline{p}_{\hat{r}(u,k)}^{ud}$ represents the probability of realization $\hat{r}(u,k)$ of random component $u$ considering the selection of distribution $d$, i.e., $\hat{p}_{\hat{r}(u,k)}^{u}(x)=\sum_{d\in D}\overline{p}_{\hat{r}(u,k)}^{ud}x_{u}^{d}$. If a unique distribution in the set $\mathcal{D}=\{1,\ldots,D\}$ is selected for the entire random vector $\boldsymbol{\xi}$ (i.e., for scenario $k\in K$), then $p_{k}(x)=\sum_{d\in\mathcal{D}}\overline{p}^{d}_{k}x^{d}$, where $\overline{p}^{d}_{k}$ is the probability of scenario $k$ according to distribution $d$, and $x^{d}$ are first-stage binary variables equal to one if distribution $d$ is selected and zero otherwise. Finally, the sum of variables $x^{d}$ over $\mathcal{D}$ must also equal one.
Dupacová (2006) also categorizes problems where decisions consist of choosing the parameters of a specific probability distribution as parameter selection. Moreover, Hellemo et al. (2018) define the probability-distortion case, where the chosen probability given $x$ is obtained by combining a set of distributions. We cannot define a general probability mass function $p_{k}(x)$ for these cases because it depends on other assumptions. Nonetheless, we characterize endogenous cases by considering the properties of the probability mass or density function of random vectors, which helps identify appropriate solution methods for each class of problems. We demonstrate this idea in Section 3.
2.2 Literature Review
Consistent with the assumptions of this work, the literature on decision-dependent probability focuses essentially on problems where the probability mass or density functions are known a priori. In this section, we review only works that use this assumption and that consider two-stage programs. We split our review into works on discrete and continuous random variables.
Discrete random variables.
For discrete random variables, the random elements are usually assumed to be independent. For the discrete selection of distributions, distinct applications of two-stage programs have been studied, such as network design protection (Viswanath et al. 2004, Flach 2010, Peeta et al. 2010, Prestwich et al. 2014), facility protection (Medal et al. 2016), facility location-protection (Li et al. 2021), fire risk management (Bhuiyan et al. 2019), supplier risk management (Zhou et al. 2022), facility protection and network design (Bhuiyan et al. 2020), and airline flight network expansion problems (Şafak et al. 2022). All these works used decision-dependent probabilities as in Equation (4), except for Şafak et al. (2022), who chose the probabilities of the entire random vector. These applications usually lead to large-scale instances with numerous scenarios.
Prestwich et al. (2014) and Bhuiyan et al. (2019) respectively obtained exact MILP reformulations of the DEP in the extensive form of the original program by using a linear scaling procedure and the probability chain technique described by O’Hanley et al. (2013). To solve larger-scale problems, Prestwich et al. (2014) proposed a bundle method, and Medal et al. (2016), Bhuiyan et al. (2020), and Zhou et al. (2022) implemented the L-shaped decomposition for solving reformulations resulting from the probability chain technique. Usually, $T_{k}$ is a zero matrix in these two-stage programs (i.e., the second stage can be solved a priori). The only exception is Zhou et al. (2022), who used an exact McCormick linearization (McCormick 1976) to obtain an MILP. These reformulations apply specifically to the case of discrete selection of distributions and rely on the binary variables in Equation (4). However, they require numerous variables and big-M constraints. Thus, the practical use of these reformulations is limited to problems with a sufficiently small number of scenarios. Flach (2010) proposed a different approximate MILP reformulation for the DEP by using the binary variables in Equation (4) to reformulate $p_{k}(x)$ as a set of exponential and linear constraints. The authors ensured a predefined precision by approximating the exponential constraints by piecewise linear cuts included in a branch-and-cut scheme.
Various decision-dependent probability mass functions have also been studied, such as the binomial distribution (Karaesmen and Van Ryzin 2004, Green et al. 2013), linear interpolation (Du and Peeta 2014), a linear scaling and convex combination of a set of distributions (Hellemo et al. 2018), and the exponential function (Bhuiyan 2021). In most studies, the endogenous uncertainty in general cases of discrete random variables resulted in DEPs, which are mixed-integer nonlinear nonconvex programs (MINLPs) for which proving global optimality is harder than for MILPs or mixed-integer convex programs (MICPs). Even if $p_{k}(x)$ are convex or concave (see, e.g., Hellemo et al. 2018) or $\hat{p}_{\hat{r}(u,k)}^{u}(x)$ are convex or concave (see, e.g., Du and Peeta 2014, Bhuiyan 2021, Holzmann and Smith 2021, Homem-de Mello et al. 2022), the term $p_{k}(x)q_{k}^{\top}y_{k}$ could lead to an MINLP. Within this context, Krasko and Rebennack (2017) studied a two-stage preparedness hazard problem for debris-flow-hazard mitigation and solved the DEP with the BARON solver. Hellemo et al. (2018) solved the DEP of an energy investment problem using a generalized Benders decomposition (GBD) and the BARON solver. However, these approaches are limited to solving instances with few scenarios because the DEP in extensive form explicitly enumerates all scenarios.
Approximative approaches have been proposed for facility protection (Medal et al. 2016), sensors for holding water quality (Karaesmen and Van Ryzin 2004), and overbooking problems (Bhuiyan 2021). The objective function in these problems is submodular, allowing the design of greedy algorithms. However, approximation guarantees are generally low (e.g., around $63\%$ for Medal et al. (2016)). Li et al. (2021) and Li et al. (2022) used genetic algorithms to solve two-stage programs of a facility location-protection problem with a bilevel program as the first stage.
Viswanath et al. (2004), Peeta et al. (2010), and Du and Peeta (2014) used the SAA within iterative heuristic procedures to evaluate current first-stage solutions for two-stage network protection problems. Unfortunately, these procedures do not provide performance guarantees. The SAA was also used to estimate the performance of current solutions within stochastic gradient methods (Karaesmen and Van Ryzin 2004, Homem-de Mello et al. 2022) for overbooking and appointment-scheduling problems, respectively. Homem-de Mello et al. (2022) proved the local convergence of the algorithm, which is compatible with its nonconvexity. Finally, by considering a specific gradient estimator, Karaesmen and Van Ryzin (2004) proved the local convergence of the algorithm.
Flach (2010) used importance sampling to apply the SAA. The author sampled from a biased probability distribution and adjusted the weight of the objective function according to $p_{k}(x)$ and the biased distribution. However, importance sampling depends on the choice of the biased distribution (Shapiro et al. 2021, Chap. 5), and finding a unique biased distribution appropriate for all conditional distributions on $x$ could be difficult. Holzmann and Smith (2021) proposed an adapted SAA for bilevel shortest-path problems with randomized strategies. They sampled realizations from the continuous uniform distribution $U(0,1)$ and included big-M constraints in the model to transform these samples into realizations of the decision-dependent discrete distribution.
Continuous random variables.
To the best of our knowledge, only Ernst and Powell (1995), Ekin et al. (2017), Ekin (2018), and Hellemo et al. (2018) investigated continuous endogenous random variables. Ernst and Powell (1995) proposed a search procedure for solving a newsvendor problem with uncertain demand following a normal distribution with mean and variance selected by first-stage decisions. The augmented probability simulation (Bielza et al. 1999), in which simulated decisions are considered to be random variables, was used by Ekin et al. (2017) and Ekin (2018) for solving, respectively, a newsvendor problem with uncertain demand depending on the definition of the stock level and a production-planning problem with random yield depending on maintenance and production decisions. In both works, the dependency was modeled with a normal distribution with mean and variance defined according to the first-stage decisions. Finally, Hellemo et al. (2018) used the discretized function of the Kumaraswamy distribution and approximated the discretized function of the normal distribution. The discretization trades an increase in the number of scenarios for a confident discrete approximation. Approximations of the discretized function may also be needed when the cumulative distribution has no closed form or has a complex closed form. In some applications, the search procedure cannot analyze the entire decision space, implying the need for a heuristic procedure. Finally, the augmented probability simulation remains limited to problems such as the newsvendor problem where the second stage is represented by analytical expressions.
3 Methodology
Our methodology is based on using transformations of random variables, which are a classical topic in probability and statistics (see Hogg et al. 1977, Chap. 5), to model endogenous uncertainty. This section outlines Assumption 3, which is necessary for applying our methodology. Section 4 shows that this assumption is not restrictive by evidencing many cases in which it holds. {assumption} There exists (1) a random vector $\vartheta^{\top}(\omega)=\boldsymbol{(}\widetilde{q}(\omega)^{\top},\widetilde{h}(\omega)^{\top},\widetilde{T}(\omega),\widetilde{W}(\omega)\boldsymbol{)}$ with a known probability distribution that is independent of the first-stage decisions $x\in\mathcal{X}$ and (2) a transformation function $t(x,\boldsymbol{\vartheta})$, such that $\boldsymbol{\xi}_{x}=t\left(x,\boldsymbol{\vartheta}\right)$. Assumption 3 allows us to mimic the endogenous uncertainty by using a function $t(\cdot,\cdot)$ that combines the first-stage variables $x$ and the decision-independent random vector $\boldsymbol{\vartheta}$. For example, $t(\cdot,\cdot)$ can be the inverse transformation of $\boldsymbol{\xi}_{x}$ (Robert and Casella 1999, Chap. 2). Here, the random vector $\boldsymbol{\vartheta}$ does not necessarily have the same dimension as $\boldsymbol{\xi}$ or $\boldsymbol{\xi}_{x}$. Furthermore, the random elements of $\boldsymbol{\xi}_{x}$ may be dependent. For the sake of simplicity, we consider that $t(x,\boldsymbol{\vartheta})=(t_{1}\boldsymbol{(}x,\widetilde{q}(\omega)^{\top}\boldsymbol{)},t_{2}\boldsymbol{(}x,\widetilde{h}(\omega)\boldsymbol{)},t_{3}\boldsymbol{(}x,\widetilde{T}(\omega)\boldsymbol{)},t_{4}\boldsymbol{(}x,\widetilde{W}(\omega)\boldsymbol{)})$, where $t_{1}\boldsymbol{(}x,\widetilde{q}\left(\omega\right)^{\top}\boldsymbol{)}$, $t_{2}\boldsymbol{(}x,\widetilde{h}\left(\omega\right)\boldsymbol{)}$, $t_{3}\boldsymbol{(}x,\widetilde{T}\left(\omega\right)\boldsymbol{)}$, and $t_{4}\boldsymbol{(}x,\widetilde{W}\left(\omega\right)\boldsymbol{)}$ are transformation functions applied to the corresponding subvectors and matrices of $\boldsymbol{\vartheta}$. These functions return, respectively, the subvectors and matrices of $\boldsymbol{\xi}_{x}$, i.e., $q_{x}(\omega)^{\top}$, $h_{x}(\omega)$, $T_{x}(\omega)$, and $W_{x}(\omega)$. From Assumption 3, $t(.,.)$ and thus $t_{1}(.,.)$, $t_{2}(.,.)$, $t_{3}(.,.)$, and $t_{4}(.,.)$, must account for any dependency between the random elements of $\boldsymbol{\xi}_{x}$. Section 4 details different methods to define appropriate transformations respecting Assumption 3.
In the field of stochastic optimization, Varaiya and Wets (1989) and Dupacová (2006) introduced the idea of transforming a decision-dependent probability distribution into an independent distribution in what is called “plug-in” or “change of measure” (see Pflug 2012 for a detailed explanation) and might employ random variable transformation. Despite these initial works, no study has directly modeled endogenous uncertainty as a function $t(.,.)$ nor investigated appropriate transformations for random variables following classical endogenous probability distributions. Specifically, the definition of $t(.,.)$ remains unexplored. Moreover, no study has focused on the nature and solution of problems resulting from this transformation nor, notably, addressed their potential nonconvexity. Finally, Holzmann and Smith (2021) used random variable transformation for a specific endogenous discrete random variable, for which Section 4.2.1 highlights the similarities and differences with the approach proposed herein. Therefore, the present work proposes a general framework using functions of random variables for addressing 2-SMIPD given Assumption 3. We also investigate practical transformations for endogenous random variables following classical probability distributions, as well as the convexity of the resulting transformed program.
Based on Assumption 3, we can formulate 2-SMIPD as
$\min$ $c^{\top}x+\mathbb{E}_{\boldsymbol{\vartheta}}\left[Q_{T}\boldsymbol{(}x,\vartheta(\omega)\boldsymbol{)}\right]$ (6a)
[…]
$t_{3}\boldsymbol{(}x,\widetilde{T}\left(\omega\right)\boldsymbol{)}x+t_{4}\boldsymbol{(}x,\widetilde{W}\left(\omega\right)\boldsymbol{)}y(\omega)\geq t_{2}\boldsymbol{(}x,\widetilde{h}\left(\omega\right)\boldsymbol{)},y(\omega)\in\mathcal{Y}\},$
and we can outline the following lemma:
Lemma 3.1
Under Assumption 3, $g(x)=\mathbb{E}_{\boldsymbol{\xi}\mid x}\left[Q\boldsymbol{(}x,\xi(\omega)\boldsymbol{)}\right]=\mathbb{E}_{\boldsymbol{\vartheta}}\left[Q_{T}\boldsymbol{(}x,\vartheta(\omega)\boldsymbol{)}\right]$ and the 2-SMIPD (original Program (1)) can be reformulated into the transformed Program (6). In other words, Programs (1) and (6) have the same optimal solutions.
Thus, our methodology is established on finding an appropriate transformation function to obtain Program (6b). In that way, under Assumption 3, 2-SMIPD can be transformed into a two-stage stochastic program with exogenous uncertainty. Solution procedures for the transformed Program (6) based on its properties are well-studied (see, e.g., Birge and Louveaux 2011, Shapiro et al. 2021). Program (6b) has the terms $t_{1}\boldsymbol{(}x,\widetilde{q}\left(\omega\right)^{\top}\boldsymbol{)}y$, $t_{2}\boldsymbol{(}x,\widetilde{h}\left(\omega\right)\boldsymbol{)}$, $t_{3}\boldsymbol{(}x,\widetilde{T}\left(\omega\right)\boldsymbol{)}x$, and $t_{4}\boldsymbol{(}x,\widetilde{W}\left(\omega\right)\boldsymbol{)}y$, which might be nonlinear nonconvex. We then introduce the following assumption: {assumption} The continuous relaxation of the second-stage Program (6b) is convex. Program (6b) might need to be reformulated to convexify its continuous relaxation. Section 4 presents different cases for which Assumption 3 holds. In addition, the exogenous random vector $\boldsymbol{\vartheta}$ might have a large or infinite dimension (see, e.g., the inverse transformation in Section 4.2). Hence, next, we describe the application of the SAA to solve Program (6), which, although intuitive, shows that the transformation does not compromise desirable SAA properties. Nonetheless, other procedures can also be employed, e.g., stochastic gradient (Birge and Louveaux 2011).
3.1 Sample Average Approximation
We can solve Program (6) with the SAA (Shapiro et al. 2021, Chap. 5). We have $N$ independent and identically distributed samples $(\boldsymbol{\vartheta^{1}},\ldots,\boldsymbol{\vartheta^{N}})$ of $\boldsymbol{\vartheta}$ with realizations $(\vartheta^{1},\ldots,\vartheta^{N})$. The SAA program is
[…]
where $\hat{g}_{N}(x)$ is the SAA estimator of $\mathbb{E}_{\boldsymbol{\vartheta}}\left[Q_{T}\boldsymbol{(}x,\vartheta(\omega)\boldsymbol{)}\right]$ defined as
$\hat{g}_{N}(x)=\frac{1}{N}\sum_{i=1}^{N}Q_{T}(x,\vartheta^{i}).$ (8)
This estimator converges point-wise with probability one (w.p.1) to $\mathbb{E}_{\boldsymbol{\vartheta}}[Q_{T}\boldsymbol{(}x,\vartheta(\omega)\boldsymbol{)}]$ when $N\rightarrow+\infty$. From Lemma 3.1, we have $\mathbb{E}[\hat{g}_{N}(x)]=g(x)$, so $\hat{g}_{N}(x)$ is an unbiased estimator of $g(x)$. We denote $v^{*}$, $v_{t(x,\boldsymbol{\vartheta})}^{*}$ and $v_{N}$ as the optimal values of Program (1), Program (6), and Program (7), respectively. Under some mild conditions, such as when Program (6b) is convex, an MILP or when the first-stage feasible set $\mathcal{X}$ is finite, $v_{N}$ converges to $v_{t(x,\boldsymbol{\vartheta})}^{*}$ w.p.1 as $N\rightarrow+\infty$ (cf. Shapiro et al. 2021, Chap. 5, Ahmed and Shapiro 2002, and Kleywegt et al. 2002, respectively). In these cases and under Assumption 3, $v_{N}$ converges to $v^{*}$ w.p.1 as $N\rightarrow+\infty$, since $v^{*}=v_{t(x,\boldsymbol{\vartheta})}^{*}$ from Lemma 3.1. This result depends on the properties of $\boldsymbol{\vartheta}$ (i.e., on the convergence of $v_{N}$ to $v_{t(x,\boldsymbol{\vartheta)}}$), which may have larger dimension and support than $\boldsymbol{\xi}$ and may require a larger number of samples.
The SAA finds a feasible first-stage solution $\overline{x}\in\mathcal{X}$. To obtain $\overline{x}$, we solve $M$ instances of the SAA Program (7), each with $N$ independent samples. We then evaluate the optimal solutions of every instance by using a larger sample size $N^{\prime}$ and define $\overline{x}$ as the optimal solution with the best evaluation. When $Q\boldsymbol{(}x,\xi(\omega)\boldsymbol{)}$ is finite and Program (1) has relatively complete recourse (i.e., the second stage associated with $x\in\mathcal{X}$ is feasible for almost every realization of $\boldsymbol{\xi}$), then $Q_{T}\boldsymbol{(}x,\vartheta(\omega)\boldsymbol{)}$ is also finite, and Program (6) also has relatively complete recourse, given that $\boldsymbol{\xi}_{x}=t(x,\boldsymbol{\vartheta})$ from Assumption 3. In this case, in addition to finding $\overline{x}$, we can compute a lower-bound estimator $\overline{v}_{M}^{N}$ of $v^{*}$, an upper-bound estimator $\hat{v}_{N^{\prime}}(\overline{x})$ of $v^{*}$, an over-estimator $\overline{gap}(\overline{x})$ of the optimality gap $c^{\top}\overline{x}+g(\overline{x})-v^{*}$, and unbiased estimators of the variances of $\overline{v}_{N}^{M}$, $\hat{g}_{N{{}^{\prime}}}(\overline{x})$, and $\overline{gap}(\overline{x})$, i.e., $\hat{\sigma}_{N,M}^{2}$, $\hat{\sigma}_{N{{}^{\prime}}}^{2}(\overline{x})$, and $\hat{\sigma}_{gap}^{2}(\overline{x})$, respectively. Appendix 8 details these estimators and the SAA’s procedure.
By defining the $\boldsymbol{\xi}_{x}$ with the function $t(x,\boldsymbol{\vartheta})$, the estimators $\hat{g}_{N}(x)$ for all $x\in\mathcal{X}$ are based on the same samples of $\boldsymbol{\vartheta}$. Choosing a transformation $t(x,\boldsymbol{\vartheta})$ that creates a positive correlation between these estimators $\hat{g}_{N}(x)$ is desirable because it reduces the variance of their difference. Indeed, when comparing two first-stage decisions $x_{1}$ and $x_{2}$, the variance of the difference $\hat{g}_{N^{\prime}}(x_{1})-\hat{g}_{N^{\prime}}(x_{2})$ when both estimators use the same random samples of $\boldsymbol{\vartheta}$ is $Var\boldsymbol{(}\hat{g}_{N{{}^{\prime}}}(x_{1})\boldsymbol{)}+Var\boldsymbol{(}\hat{g}_{N{{}^{\prime}}}(x_{2})\boldsymbol{)}-2Cov\boldsymbol{(}\hat{g}_{N{{}^{\prime}}}(x_{1}),\hat{g}_{N{{}^{\prime}}}(x_{2})\boldsymbol{)}$, where $Var(\cdot)$ and $Cov(\cdot,\cdot)$ denote the variance and covariance. This is the reasoning behind the well-known efficiency-improvement technique of common random numbers (Glasserman and Yao 1992). This discussion extends to the application of any sampling technique for solving the transformed program. Section 4 further discusses this question.
4 Transformation Functions: Definition and Examples
This section presents appropriate transformations $\boldsymbol{\xi}_{x}=t(x,\boldsymbol{\vartheta})$ with $x\in\mathcal{X}$ for classical distributions in which both Assumptions 3 and 3 hold. We classify the transformation approaches according to four main categories to organize our discussion: enumeration (Section 4.1), inversion (Section 4.2), standardization (Section 4.3), and convolution (Section 4.4). Then, Section 4.5 discusses the proposed methodology’s advantages and disadvantages. We use the notation introduced in Section 2.1, and we denote $\boldsymbol{\xi}_{x}^{u}$ as the $u$-th element of the random vector $\boldsymbol{\xi}_{x}$, and $\boldsymbol{\vartheta}_{u}$ as the random variable or vector used to obtain $\boldsymbol{\xi}_{x}^{u}$. Similarly, $\xi_{x}^{u}$ and $\vartheta_{u}$ denote their realizations. Unless stated otherwise, we consider the random elements of $\boldsymbol{\xi}_{x}$ to be independent. If the random elements of $\boldsymbol{\xi}_{x}$ are independent, then the $\boldsymbol{\vartheta}_{u}$ for $u\in\mathcal{U}$ must also be independent. Thus, we can model each random element $\boldsymbol{\xi}_{x}^{u}$ separately by using its marginal probability distribution conditioned on $x$ and the $\boldsymbol{\vartheta}_{u}$. Nonetheless, in certain cases, when the elements of $\boldsymbol{\xi}_{x}$ are correlated, we can have correlated $\boldsymbol{\vartheta}_{u}$ and also model each $\boldsymbol{\xi}_{x}^{u}$ separately. As a result, we mainly focus on the individual transformation of an element $u$ in this section, i.e., $\boldsymbol{\xi}_{x}^{u}=t(x,\boldsymbol{\vartheta}_{u})$. We remark that most studies have focused on the independent case (see Section (2.2)), while our approach is suitable for both independent and dependent cases.
4.1 Enumeration
For a random element $\boldsymbol{\xi}_{x}^{u}$ with $u\in\mathcal{U}$, a simple transformation can be employed if the set of alternative marginal distributions $\mathcal{D}_{u}$ of element $u$ is sufficiently small to allow enumeration. This corresponds to the “discrete selection of distributions” (see Section 2.1). We have $\boldsymbol{\vartheta}_{u}=\{\boldsymbol{\vartheta}_{ud}\}_{d\in\mathcal{D}_{u}}$, where $\boldsymbol{\vartheta}_{ud}$ is a random variable following distribution $d\in\mathcal{D}_{u}$. The $\boldsymbol{\vartheta}_{u}$ is independent of $x$ and respects Property 1 of Assumption 3. We thus have $t(x,\boldsymbol{\vartheta}_{u})=\sum_{d=1}^{D_{u}}\boldsymbol{\vartheta}_{ud}x_{u}^{d}$, with $x_{u}^{d}$ as defined in Section 2.1, and Property 2 of Assumption 3 holds. The elements of $\boldsymbol{\vartheta}_{u}$ may be independent or correlated (e.g., by using common random numbers). We exemplify this in Example 4.1.
Example 4.1
Consider two independent random elements $\boldsymbol{\xi}_{x}^{A}$ and $\boldsymbol{\xi}_{x}^{B}$, with set of realizations $\mathcal{R}_{A}=\{r_{A}^{1},r_{A}^{2},r_{A}^{3}\}$ and $\mathcal{R}_{B}=\{r_{B}^{1},r_{B}^{2},r_{B}^{3}\}$, respectively. Thus, the original program has nine scenarios. These random variables have a set of alternative marginal distributions $\mathcal{D}_{A}=\{d_{A}^{1},d_{A}^{2}\}$ and $\mathcal{D}_{B}=\{d_{B}^{1},d_{B}^{2}\}$, respectively. Only realizations $r_{A}^{1}$ and $r_{A}^{2}$ have a probability greater than zero in distribution $d_{A}^{1}$, while all three realizations have a probability greater than zero in distribution $d_{A}^{2}$ (similarly for $d_{B}^{1}$ and $d_{B}^{2}$). If $d_{A}^{1}$ and $d_{B}^{1}$ are chosen, then only four scenarios have a probability greater than zero. However, if $d_{A}^{2}$ and $d_{B}^{2}$ are chosen, all nine scenarios have positive probabilities. We have two independent exogenous random vectors: $\boldsymbol{\vartheta}_{A}=\{\boldsymbol{\vartheta}_{A,d_{A}^{1}},\boldsymbol{\vartheta}_{A,d_{A}^{2}}\}$ and $\boldsymbol{\vartheta}_{B}=\{\boldsymbol{\vartheta}_{B,d_{B}^{1}},\boldsymbol{\vartheta}_{B,d_{B}^{2}}\}$. While $\boldsymbol{\vartheta}_{A}$ and $\boldsymbol{\vartheta}_{B}$ are independent, their elements can be independent or correlated, i.e., $\boldsymbol{\vartheta}_{A,d_{A}^{1}}$ can be correlated with $\boldsymbol{\vartheta}_{A,d_{A}^{2}}$, similarly for $\boldsymbol{\vartheta}_{B,d_{B}^{1}}$ and $\boldsymbol{\vartheta}_{B,d_{B}^{2}}$. The random variable $\boldsymbol{\vartheta}_{A,d_{A}^{1}}$ follows distribution $d_{A}^{1}$ and has two possible realizations ($r_{A}^{1}$ and $r_{A}^{2}$), while $\boldsymbol{\vartheta}_{A,d_{A}^{2}}$ follows distribution $d_{A}^{2}$ and has three possible realizations ($r_{A}^{1}$, $r_{A}^{2}$ and $r_{A}^{3}$). The reasoning is the same for $\boldsymbol{\vartheta}_{B,d_{B}^{1}}$ and $\boldsymbol{\vartheta}_{B,d_{B}^{2}}$. Hence, the transformed program has 36 scenarios. If distributions $d_{A}^{1}$ and $d_{B}^{1}$ are chosen, then $t(x,\boldsymbol{\vartheta}_{A})$ and $t(x,\boldsymbol{\vartheta}_{B})$ return random variables $\boldsymbol{\vartheta}_{A,d_{A}^{1}}$ and $\boldsymbol{\vartheta}_{B,d_{B}^{1}}$, respectively. Thus, for the scenario composed by realizations $r_{A}^{1}$, $r_{A}^{3}$, $r_{B}^{1}$, $r_{B}^{3}$ for random variables $\boldsymbol{\vartheta}_{A,d^{1}_{A}}$, $\boldsymbol{\vartheta}_{A,d^{2}_{A}}$, $\boldsymbol{\vartheta}_{B,d^{1}_{B}}$ and $\boldsymbol{\vartheta}_{B,d^{2}_{B}}$, respectively, the functions $t(x,\boldsymbol{\vartheta}_{A})$ and $t(x,\boldsymbol{\vartheta}_{B})$ return realizations $r_{A}^{1}$ and $r_{B}^{1}$.
If an enumeration of the alternative distributions $\mathcal{D}$ of the entire random vector $\boldsymbol{\xi}$ is viable, then we can apply the same transformation considering all the elements of $\boldsymbol{\xi}$. That is, $t(x,\boldsymbol{\vartheta})=\sum_{d=1}^{D}\boldsymbol{\vartheta}_{d}x^{d}$, where $\boldsymbol{\vartheta}_{d}$ follows distribution $d\in\mathcal{D}$ and $x^{d}$ is defined as in Section 2.1. This transformation can be applied for $\boldsymbol{\xi}_{x}$ with independent or dependent random elements. Finally, if bilinear terms appear due to $t_{1}\boldsymbol{(}x,\widetilde{q}(\omega)^{\top}\boldsymbol{)}y$, $t_{3}\boldsymbol{(}x,\widetilde{T}(\omega)\boldsymbol{)}x$ and $t_{4}\boldsymbol{(}x,\widetilde{W}(\omega)\boldsymbol{)}y$, then auxiliary linearizations such as McCormick linearization (McCormick 1976) can be applied so Assumption 3 is respected.
4.2 Inversion
The inversion method is commonly used to sample a random variable when the probability mass or density function is known (e.g., the inverse sampling method, cf. Robert and Casella 1999, Chap. 2). This technique transforms a random variable following the continuous uniform distribution $U(0,1)$ into the desired random variable. We define $\mathbb{F}_{\boldsymbol{\xi}^{u}}(x,\beta)(\cdot)$ to be the conditional marginal cumulative distribution function of the random variable $\boldsymbol{\xi}^{u}$ given $x$, the first-stage variables, and $\beta$, a set of remaining parameters that are not decisions of the optimization model. To apply the inverse transformation, we compute the generalized inverse of the conditional cumulative distribution:
$\mathbb{F}^{-1}_{\boldsymbol{\xi}^{u}}(x,\beta)(j)=\min\left\{\xi\in\Xi_{x}\mid\mathbb{F}_{\boldsymbol{\xi}^{u}}(x,\beta)(\xi)\geq j\right\}.$ (9)
Hence, the transformation is defined as $\boldsymbol{\xi}_{x}^{u}=t(x,\boldsymbol{\vartheta}_{u})=\mathbb{F}^{-1}_{\boldsymbol{\xi}^{u}}(x,\beta)(\boldsymbol{\vartheta}_{u})$ (Property 2 of Assumption 3), such that $\boldsymbol{\vartheta}_{u}$ is a random variable following $U(0,1)$ and independent of $x$ (Property 1 of Assumption 3). The inversion method, in many cases, has the desirable property of inducing a positive correlation (see Section 3.1) between the estimators of different systems (i.e., the $\hat{g}_{N}(x)$ for $x\in\mathcal{X}$) when their samples are generated from the same sequence of uniform random numbers (Whitt 1976).
If the elements of $\boldsymbol{\xi}_{x}$ are dependent and their correlation matrix is known, then we can define correlated random elements $\boldsymbol{\vartheta}_{u}$ for $u\in\mathcal{U}$ using an appropriate copula. A copula is a multivariate cumulative distribution function that ensures the marginal distribution of each random variable $\boldsymbol{\vartheta}_{u}$ is the $U(0,1)$. Copulas are well-known for modeling dependencies (see, e.g., Nelsen 2006). Thus, as in the independent case, each $\boldsymbol{\vartheta}_{u}$ can be transformed into $\boldsymbol{\xi}_{x}^{u}$ separately using the marginal distribution of this endogenous random element. Example 4.2 illustrates the application of the inverse.
Example 4.2
Consider two independent (correlated) random variables $\boldsymbol{\xi}_{x}^{A}$ and $\boldsymbol{\xi}_{x}^{B}$ following the marginal cumulative distribution $\mathbb{F}_{\boldsymbol{\xi}^{A}}(x,\beta)(\cdot)$ and $\mathbb{F}_{\boldsymbol{\xi}^{B}}(x,\beta)(\cdot)$, respectively. We have two independent (correlated) random variables $\boldsymbol{\vartheta}_{A}$ and $\boldsymbol{\vartheta}_{B}$ following the marginal distribution $U(0,1)$. For the independent case, the $\boldsymbol{\vartheta}_{A}$ and $\boldsymbol{\vartheta}_{B}$ follow the independence bi-variate copula. For the correlated case, the $\boldsymbol{\vartheta}_{A}$ and $\boldsymbol{\vartheta}_{B}$ follow a bi-variate copula which models the correlation of $\boldsymbol{\xi}_{x}^{A}$ and $\boldsymbol{\xi}_{x}^{B}$, e.g., the Gaussian copula with the corresponding correlation. A scenario is composed of the realizations of $\boldsymbol{\vartheta}_{A}$ and $\boldsymbol{\vartheta}_{B}$, which are transformed into $\boldsymbol{\xi}_{x}^{A}$ and $\boldsymbol{\xi}_{x}^{B}$ by their respective transformation $t(x,\boldsymbol{\vartheta}_{u})=\mathbb{F}^{-1}_{\boldsymbol{\xi}^{u}}(x,\beta)$ for $u\in\{A,B\}$.
Next, we investigate the application of this technique for discrete random variables and continuous random variables according to the properties of their cumulative probability distribution.
4.2.1 Discrete Random Variables.
In this case, the inverse transformation $\boldsymbol{\xi}_{x}^{u}=\mathbb{F}^{-1}_{\xi^{u}}(x,\beta)(\boldsymbol{\vartheta}_{u})$ is simplified and consists of defining the realization $r\in\mathcal{R}_{u}$ such that
$\sum_{j=1}^{r-1}\hat{p}_{j}^{u}(x)\leq\vartheta_{u}<\sum_{j=1}^{r}\hat{p}_{j}^{u}(x),$ (10)
where the sum on the left-hand-side equals zero when $r=1$, and $\hat{p}_{j}^{u}(x)$ and $\sum_{j=1}^{r}\hat{p}_{j}^{u}(x)$ are the marginal and cumulative probability functions (see Section 2.1) of the realization $r$ of random element $u\in\mathcal{U}$. Thus, we define the variables $\theta_{u}^{r}\in\{0,1\}$, which take the value one if the realization $r\in\mathcal{R}_{u}$ respects Inequalities (10). We also have the variables $\pi_{u}^{r}\in\{0,1\}$, which equal the sum $\sum_{j=1}^{r}\theta_{u}^{j}$ for every realization $r\in\mathcal{R}_{u}$. The following constraints model Inequalities (10):
$\pi_{u}^{r}\leq\frac{1}{\vartheta_{u}}\sum_{j=1}^{r}\hat{p}_{j}^{u}(x)-\epsilon$ $\forall\ r\in\{1,\ldots,R_{u}-1\}$ (11)
$(1-\pi_{u}^{r-1})\leq\frac{1}{(1-\vartheta_{u})}\sum_{j=r}^{R_{u}}\hat{p}_{j}^{u}(x)$ $\forall\ r\in\{2,\ldots,R_{u}\}$ (12)
$\pi_{u}^{r}=\pi_{u}^{r+1}-\theta_{u}^{r+1}$ $\forall\ r\in\{1,\ldots,R_{u}-1\}$ (13)
$\sum_{r\in\mathcal{R}_{u}}\theta_{u}^{r}=1,$ (14)
where $\epsilon>0$ is a sufficiently small value used to define the strict inequality in (10). Constraints (11) ensure that the right-hand side in Inequalities (10) is respected. Constraints (12) certify that the left-hand side in Inequalities (10) is respected. Constraints (13) establish the connection between $\pi^{r}_{u}$ and $\theta^{r}_{u}$ so that $\theta^{r}_{u}=1$ for the right realization, where $\pi_{u}^{R_{u}}$ is set to one. Although we could directly model Constraints (11) and (12) with variables $\theta_{u}^{r}$, we use the variables $\pi_{u}^{r}$ to reduce the number of nonzero elements in these constraints. Constraint (14) restricts the choice of a unique realization for the random element $u$. Finally, $\nu_{r}^{u}$ is the value of realization $r$ of the random element $u$. Thus, $t(x,\vartheta_{u})=\sum_{r=1}^{R_{u}}\nu_{r}^{u}\theta_{u}^{r}$. To ensure Assumption 3, the continuous relaxation of Constraints (11) and (12) must be convex. This leads to the following theorem:
Theorem 4.3
If the functions $\hat{p}^{u}_{r}(x)$ for $u\in\mathcal{U}$ and $r\in\mathcal{R}_{u}$ have a concave continuous relaxation on $x\in\mathcal{X}$, Constraints (11) and (12) are convex. The same is true if these functions have a convex continuous relaxation.
Appendix 9 presents the proof of Theorem 4.3. In addition, even if the continuous relaxation of the $\hat{p}_{r}^{u}(x)$ is neither concave nor convex, we can convexify Constraints (11) and (12) in certain cases. This analysis is problem-specific. Different studies investigate problems with convex or concave $p_{r}^{u}(x)$ (see Section 2.2). Despite this desirable property, the original program is usually nonconvex. In contrast, using the inverse transformation function results in only linear or convex constraints.
To the best of our knowledge, only Holzmann and Smith (2021) use the discrete inverse transformation to transform the endogenous uncertainty in 2-SMIPD into exogenous uncertainty for a specific problem. In contrast, we describe a general framework adapted to various functions $\hat{p}_{r}^{u}(x)$. Moreover, the set of Constraints (11)–(14) has a tighter continuous relaxation feasibility set than that used by Holzmann and Smith (2021) (see Appendix 10). Finally, Holzmann and Smith (2021) do not investigate the statistical estimators obtained with their formulation, which we study for general transformations (Section 3) and also validate with a case study in Sections 5 and 6.
In certain cases, such as when $\boldsymbol{\xi}_{x}^{u}$ has only two possible realizations and $p_{r}^{u}(x)$ is a linear function, simpler models can be designed. This is the case of the Bernoulli distribution, described next.
Bernoulli Distribution.
We consider that $\boldsymbol{\xi}_{x}^{u}\sim Bern(\phi)$, with the probability of success $\phi\in[0,1]$ being a subvector of $x$. The inverse is such that $\xi_{x}^{u}=1$ if $\vartheta_{u}<\phi$, and $\xi_{x}^{u}=0$ otherwise, with $\boldsymbol{\vartheta}_{u}\sim U(0,1)$. We model this with the variables $\upsilon_{u}\in\{0,1\}$, i.e., $t(x,\vartheta_{u})=\upsilon_{u}$, and the constraints
$\{\vartheta_{u}\upsilon_{u}\leq\phi-\epsilon\;;\;\upsilon_{u}\geq\phi-\vartheta_{u}\},$ (15)
where $\epsilon>0$ is again a sufficiently small value. Constraints (15) guarantee that realization 00 and $1$ are chosen if $\vartheta_{u}$ is superior to $\phi$ and if $\vartheta_{u}$ is inferior or equal to $\phi$, respectively.
For both transformations (general and specific to Bernoulli), linearization techniques, such as McCormick, can be easily applied to linearize the bilinear terms in $t_{1}\boldsymbol{(}x,\widetilde{q}(\omega)^{\top}\boldsymbol{)}y$, $t_{3}\boldsymbol{(}x,\widetilde{T}(\omega)\boldsymbol{)}x$, and $t_{4}\boldsymbol{(}x,\widetilde{W}(\omega)\boldsymbol{)}y$, as they involve at least one binary variable. Thus, Assumption 3 holds.
4.2.2 Continuous Random Variables.
In general, the transformation $\xi_{x}^{u}=\mathbb{F}^{-1}_{\xi^{u}}(x,\beta)(\vartheta_{u})$ cannot be simplified for continuous random variables. Nevertheless, when, for fixed $x$ and $\beta$, the conditional cumulative distribution is strictly increasing and continuous, the generalized inverse is redundant because $\mathbb{F}^{-1}_{\boldsymbol{\xi}^{u}}(x,\beta)(\cdot)$ is the unique inverse function. This property is commonly respected for well-behaved distributions and leads to a simpler transformation. Two applications of the inverse for some well-known continuous distributions respecting this property are listed below.
Uniform distribution.
If $\boldsymbol{\xi}^{u}\sim U(a,b)$ and both parameters $a\in\mathbb{R}$ and $b\in\mathbb{R}$ are first-stage variables, then we can use the transformation function $t(a,b,\boldsymbol{\vartheta}_{u})=a+(b-a)\boldsymbol{\vartheta}_{u}$.
Exponential distribution.
If $\boldsymbol{\xi}^{u}\sim\exp(\lambda)$ with $\lambda>0$ a first-stage decision, then the transformation is defined as $t(\lambda,\boldsymbol{\vartheta}_{u})=-{\ln(1-\boldsymbol{\vartheta}_{u})}/{\lambda}$. This function is concave for $\lambda>0$.
Both distributions are part of the special case where the unique inverse $\mathbb{F}^{-1}_{\boldsymbol{\xi}^{u}}(x,\beta)(\cdot)$ is convex or concave on $x\in\mathcal{X}$, i.e., the inverse function is convex or concave on the parameters of the distribution. Other distributions that fall into this special case are the Cauchy and logistic distributions. Additionally, there are distributions for which $\mathbb{F}^{-1}_{\boldsymbol{\xi}^{u}}(x,\beta)(\cdot)$ has no closed-form expression, but good numerical approximations are available, e.g. student, chi-square, and normal distributions. Nevertheless, transformations based on other techniques, which lead to simpler formulations, can be used for certain distributions (see Section 4.3 for the normal distribution). The analysis of the special class of distributions introduced before leads to the following theorem:
Theorem 4.4
Assumption 3 holds if the following criteria are met: (1) the cost vector $q$, technology matrix $T$ and recourse matrix $W$ are fixed; (2) $\mathbb{F}_{\boldsymbol{\xi}^{u}}(x,\beta)(\cdot)$for fixed $x$ and $\beta$ is continuous and strictly increasing; and (3) the transformation $t(x,\boldsymbol{\vartheta}_{u})=\mathbb{F}^{-1}_{\boldsymbol{\xi}^{u}}(x,\beta)(\cdot)$ is a convex function on $x\in\mathcal{X}$, or concave if the recourse constraint in Program (6b) is an inferior or equal to inequality.
Appendix 9 presents the proof of Theorem 4.4. If Criterion (1) does not hold, the products $t_{1}\boldsymbol{(}x,\widetilde{q}(\omega)^{\top}\boldsymbol{)}y$, $t_{3}\boldsymbol{(}x,\widetilde{T}(\omega)\boldsymbol{)}x$ and $t_{4}\boldsymbol{(}x,\widetilde{W}(\omega)\boldsymbol{)}y$ may be nonconvex even if Criterion (3) holds. Thus, a problem-specific convexity analysis is required to confirm if Assumption 3 is satisfied.
4.3 Standardization
A location-scale family of distributions has distributions with location $\mu$ and scale $\sigma$ parameters (Casella and Berger 2024, Chap. 3.5). The distributions of this family can be derived from a standard family distribution member with location and scale parameters equal to one and zero, respectively. If $\boldsymbol{\xi}^{u}$ follows a location-scale family distribution with parameters $\mu$ and $\sigma$ as first-stage decisions, then $t(\mu,\sigma,\boldsymbol{\vartheta}_{u})=\mu+\sigma\boldsymbol{\vartheta}_{u}$ is valid (Property 2 of Assumption 3). The $\boldsymbol{\vartheta}_{u}$ follows the standard family distribution member, which is independent of the values of $\mu$ and $\sigma$ (Property 1 of Assumption 3). If only the vector $h(\omega)$ has random elements, then Assumption 3 holds. Otherwise, a specific analysis of the terms $t_{1}(x,\widetilde{q}(\omega)^{\top})y$, $t_{3}(x,\widetilde{T}(\omega))x$ and $t_{4}(x,\widetilde{W}(\omega))y$ should be conducted. Many distributions have this property, such as the normal, generalized extreme value, exponential, and logistic distributions. Next, we exemplify this method for two of these distributions.
Normal distribution.
If $\boldsymbol{\xi}^{u}\sim N(\mu,\sigma^{2})$ and $\mu$ and $\sigma$ are first-stage decisions, then we can apply the proposed transformation with $\boldsymbol{\vartheta}_{u}$ following the standard normal $N(0,1)$.
Generalized extreme value distribution.
If $\boldsymbol{\xi}^{u}\sim GEV(\mu,\sigma,\beta)$ with $\mu$ and $\sigma$ first-stage decisions and $\beta$ a fixed parameter, then we can apply the proposed transformation with $\boldsymbol{\vartheta}_{u}\sim GEV(0,1,\beta)$.
4.4 Convolution
Various random variables that follow well-known distributions can be expressed as the sum of other independent random variables. Notable examples include the binomial distribution, which is the sum of Bernoulli random variables; and the hypoexponential distribution, which is the sum of exponential random variables. Suppose that $\boldsymbol{\xi}_{x}^{u}$ is equal to the sum $\boldsymbol{\zeta}_{x}^{u1}+\boldsymbol{\zeta}_{x}^{u2}+\ldots+\boldsymbol{\zeta}_{x}^{un}$ of independent random variables $\boldsymbol{\zeta}_{x}^{uj}$’s following a specified endogenous distribution. In this case, the probability distribution of $\boldsymbol{\xi}_{x}^{u}$ is the convolution of the marginal distributions of the $\boldsymbol{\zeta}_{x}^{uj}$’s  (Hogg et al. 2013, Chap. 2). We can define a transformation $\boldsymbol{\zeta}_{x}^{uj}=\hat{t}_{j}(x,\boldsymbol{\vartheta}_{uj})$ for every endogenous random variable $\boldsymbol{\zeta}_{x}^{uj}$, $j\in\{1,\ldots,n\}$, independently. Thus, $t(x,\boldsymbol{\vartheta}_{u})=\sum_{j=1}^{n}\hat{t}_{j}(x,\boldsymbol{\vartheta}_{uj})$, with $\boldsymbol{\vartheta}_{u}=\{\boldsymbol{\vartheta}_{uj}\}_{j=1}^{n}$. Next, we demonstrate this transformation for the binomial distribution. We also provide an example of dependent random variables for which this methodology can be applied.
Binomial distribution.
We consider that $\boldsymbol{\xi}_{x}^{u}\sim Bin(n,\phi)$, where the probability of success of one trial $\phi\in[0,1]$ is a subvector of $x$, and the number of trials $n$ is a parameter. We have $\boldsymbol{\xi}_{x}^{u}=\boldsymbol{\zeta}_{x}^{u1}+\boldsymbol{\zeta}_{x}^{u2}+\ldots+\boldsymbol{\zeta}_{x}^{un}$, where the $\boldsymbol{\zeta}_{x}^{uj}$’s are endogenous and follow the $Bern(\phi)$ independently. We define the $\boldsymbol{\zeta}_{x}^{uj}$’s with the inversion for the Bernoulli distribution of Section 4.2.1. We have $\boldsymbol{\vartheta}_{uj}\sim U(0,1)$ independently for $j\in\{1,\ldots,n\}$. We set the variables $\upsilon_{ur}\in\{0,1\}$ equal to one if realization $\vartheta_{ur}<\phi$ and zero otherwise, and include Constraints (15) for all $\boldsymbol{\zeta}_{x}^{uj}$’s. Thus, $t(x,\vartheta_{u})=\sum_{j=1}^{n}\upsilon_{uj}$. The bilinear terms in $t_{1}(x,\widetilde{q}(\omega)^{\top})y$, $t_{3}(x,\widetilde{T}(\omega))x$, and $t_{4}(x,\widetilde{W}(\omega))y$ can be linearized by using the McCormick envelope (Assumption 3 holds). The inverse transformation (see Section 4.2.1) may be used in this case because the binomial probability of every realization is a convex function on the probability of success $\phi$ (i.e., we can apply Theorem 4.3). Both approaches have a similar number of variables and constraints and include big-M constraints. However, the general approach leads to convex (potentially nonlinear) constraints, whereas convolution leads to linear constraints.
Example 4.5
Consider two correlated random elements $\boldsymbol{\xi}_{x}^{A}$ and $\boldsymbol{\xi}_{x}^{B}$, such that $\boldsymbol{\xi}_{x}^{A}=\boldsymbol{\zeta}_{x}^{A1}+\boldsymbol{\zeta}_{x}^{AB2}$ and $\boldsymbol{\xi}_{x}^{B}=\boldsymbol{\zeta}_{x}^{B1}+\boldsymbol{\zeta}_{x}^{AB2}$. In this case, the correlation is a result of the dependency on $\boldsymbol{\zeta}_{x}^{AB2}$. Therefore, if $\boldsymbol{\zeta}_{x}^{AB2}=\hat{t}_{AB2}(x,\boldsymbol{\vartheta}_{AB2})$, $\boldsymbol{\zeta}_{x}^{A1}=\hat{t}_{A1}(x,\boldsymbol{\vartheta}_{A1})$, and $\boldsymbol{\zeta}_{x}^{B1}=\hat{t}_{B1}(x,\boldsymbol{\vartheta}_{B1})$, then $\boldsymbol{\xi}_{x}^{A}=\hat{t}(x,\boldsymbol{\vartheta}_{A1})+\hat{t}(x,\boldsymbol{\vartheta}_{AB2})$ and $\boldsymbol{\xi}_{x}^{B}=\hat{t}(x,\boldsymbol{\vartheta}_{B1})+\hat{t}(x,\boldsymbol{\vartheta}_{AB2})$, with $\boldsymbol{\vartheta_{A}}=\{\boldsymbol{\vartheta}_{A1},\boldsymbol{\vartheta}_{AB2}\}$ and $\boldsymbol{\vartheta_{B}}=\{\boldsymbol{\vartheta}_{B1},\boldsymbol{\vartheta}_{AB2}\}$.
4.5 Advantages and Disadvantages
The proposed method obtains the transformed Program (6), a two-stage stochastic program with exogenous uncertainty. This transformed program offers several advantages: (i) Solution procedures for such programs are extensively researched and can be applied based on their specific properties. (ii) Large-scale decomposition techniques, such as the L-shaped algorithm, are applicable depending on the characteristics of the transformed second-stage Program (6b). Besides, this method is versatile as various techniques can be employed to define the transformation function $t(x,\boldsymbol{\vartheta})$. As a result, multiple transformations may exist to model the same endogenous random vector. For instance, both the inverse and convolution methods can be applied to model the binomial random variable. The choice of a transformation should aim to achieve a transformed program with desirable properties. Furthermore, as demonstrated earlier in this section, in many cases, a transformation function leading to a mixed-integer linear or convex transformed Program (6) exists, even if the original Program (1) is nonlinear nonconvex. Even if the original program with endogenous uncertainty is an MILP, a simpler transformed program can still be obtained. For the discrete selection of distributions, the original program is an MILP obtained by linearizing the objective function with many variables and big-M constraints (see (Bhuiyan et al. 2020)). However, we propose a transformation based on the enumeration of distributions (see Section 4.1) that reduces considerably the need for additional variables or constraints.
Nevertheless, certain transformations, such as the inverse transformation for discrete random variables, may introduce many variables and constraints in the formulation. Hence, we explore the impact of such transformations in our experimental results. Additionally, the dimension and support of the exogenous random vector $\boldsymbol{\vartheta}$ are typically larger than those of the endogenous random vector $\boldsymbol{\xi}_{x}$. The convergence properties of sampling techniques depend on the support of $\boldsymbol{\vartheta}$, potentially requiring more samples. However, leveraging the same random samples of $\boldsymbol{\vartheta}$ for the SAA estimators $\hat{g}_{N}(x)$ for all $x\in\mathcal{X}$ to induce correlation between them is consistent with the efficiency-improvement technique of common random numbers (see discussion in Section 3.1).
5 Case Study
Pre-disaster or pre-interdiction problems with endogenous uncertainty have been investigated in the context of facility protection (Medal et al. 2016, Bhuiyan et al. 2020, Li et al. 2021, 2022) and network design protection (Viswanath et al. 2004, Flach 2010, Peeta et al. 2010, Du and Peeta 2014, Prestwich et al. 2014). To the best of our knowledge, these applications are the most recurrent in previous research. Thus, this section showcases our methodology by applying it to the NDFPP, which was introduced by Bhuiyan et al. (2020).
Bhuiyan et al. (2020) defined the endogenous random variables through a discrete selection of distributions. In contrast, we explore this and three other decision-dependent probability mass or density functions for the random variables. This section introduces our transformed program for the NDFPP, which has different versions according to the definition of the endogenous distribution.
5.1 Network Design and Facility Protection Problem
The NDFPP is a 2-SMIPD with complete recourse. First-stage decisions determine the allocation of protections for facilities and the network structure. Next, the scenario (natural disruptions) is realized, affecting the facilities’ capacity levels, and the second-stage decisions establish an optimal flow to satisfy client demands. The capacity levels of facilities are mutually independent endogenous random variables.
We consider a set $\mathcal{D}_{0}$ of natural disruption events, which are different types of natural disasters with a center of occurrence. Facilities are exposed to distinct natural disruption intensity levels depending on their distance to the center of disruption $d\in\mathcal{D}_{0}$. The disruption intensity levels populate the set =$\mathcal{L}=\{1,\ldots,L\}$. We define the set $\mathcal{D}=\mathcal{D}_{0}\cup\{nd\}$ of possible events, where $nd$ is the no-natural-disruption event. The network is represented by the undirected graph $G=(\mathcal{N}_{\mathcal{T}},\mathcal{E})$, where $\mathcal{N}_{\mathcal{T}}=\mathcal{N_{F}}\cup\mathcal{N_{C}}\cup\{m\}$, with $\mathcal{N_{F}}$, $\mathcal{N_{C}}$, and $m$ being the set of facilities, the set of clients, and a dummy facility, respectively. In addition, $\mathcal{N}_{\mathcal{T}}^{0}=\mathcal{N}_{\mathcal{T}}\backslash\{m\}$ is the set of nodes without the dummy facility. We also specify a directed graph $G_{o}=(\mathcal{N}_{\mathcal{T}},\mathcal{A})$ by including the arcs $(i,j)$ and $(j,i)$ in $\mathcal{A}$ for each edge $[i,j]\in\mathcal{E}$. We denote as $e(i,j)$ the edge $e\in\mathcal{E}$ corresponding to arc $(i,j)\in\mathcal{A}$. Moreover, to represent the unsatisfied demand, we include arcs $(m,n)$ for every node $n\in\mathcal{N_{C}}$ in the set $\mathcal{A}$. Thus, the auxiliary set of arcs arriving at the dummy facility is defined as $\mathcal{A}_{m}=\{(m,n)\ \forall\ n\in\mathcal{N_{C}}\}$. The set $\mathcal{P}=\{1,\ldots,P\}$ contains the available protection levels for each facility, where larger values represent better protection. Finally, $\mathcal{S}$ is the set of scenarios, where each scenario is composed of an event and a realization of the exogenous random vector. We define the function $\hat{d}(s)$, which returns the event $d\in\mathcal{D}$ of scenario $s\in\mathcal{S}$. The realization of the exogenous random vector in scenario $s\in\mathcal{S}$ is denoted $\vartheta^{s}$. The definition of this random vector depends on the endogenous distribution and transformation function, so we define it later.
Protecting a facility $f\in\mathcal{N_{F}}$ with protection level $p\in\mathcal{P}$ has cost $c_{f}^{p}$ and opening edge $e\in\mathcal{E}$ has cost $c_{e}$. Every client node $n\in\mathcal{N_{C}}$ has demand $b_{n}$. We denote as $C$ the budget available for edge opening and facility protection. The per-unit-flow transportation cost of arc $(i,j)\in\mathcal{A}$ is defined as $q_{ij}$. The cost of arcs $\mathcal{A}_{m}$ leaving the dummy facility is $a\cdot\max_{(i,j)\in\mathcal{A}\backslash\mathcal{A}_{m}}q_{ij}$, where $a$ is the penalty per unit of unmet demand. Finally, $B=\sum_{n\in\mathcal{N_{C}}}b_{n}$ is the big-M value for edge capacity.
For the first stage, we define the binary variables $x_{f}^{p}$, which are equal to one if the protection level $p\in\mathcal{P}$ is allocated to facility $f\in\mathcal{N_{F}}$ and zero otherwise. We also include the binary variables $z_{e}$, equal to one if edge $e\in\mathcal{E}$ is installed and zero otherwise. For the second stage, we specify the variables $y_{ij}^{s}$ as the flow passing through arc $(i,j)\in\mathcal{A}$ in scenario $s\in\mathcal{S}$. The endogenous capacity of the facilities depends on the $x$ variables. Thus, following the notation of Section 4, we denote as $\xi_{x}^{fs}$ the endogenous capacity of facility $f\in\mathcal{N_{F}}$ in scenario $s\in\mathcal{S}$. We summarize this notation in Appendix 14. The general transformed program for the NDFPP is
$\min\mbox{: }$ $\mathbb{E}_{\boldsymbol{\vartheta}}[Q_{T}(x,\boldsymbol{\vartheta})]$ (16a)
s.t. $\sum_{p\in\mathcal{P}}x_{f}^{p}=1$ $\quad\forall\ f\in\mathcal{N_{F}}$ (16b)
$\sum_{f\in\mathcal{N_{F}}}\sum_{p\in\mathcal{P}}c_{f}^{p}x_{f}^{p}+\sum_{e\in\mathcal{E}}c_{e}z_{e}\leq C$ (16c)
$x_{f}^{p}\in\{0,1\}$ $\quad\forall\ f\in\mathcal{N_{F}},\quad p\in\mathcal{P}$ (16d)
$z_{e}\in\{0,1\}$ $\quad\forall\ e\in\mathcal{E},$ (16e)
where the transformed second-stage program $Q_{T}(x,\vartheta_{s})$ for realization $\vartheta_{s}$ of $\boldsymbol{\vartheta}$ is
$\min\mbox{: }$ $\sum_{(i,j)\in\mathcal{A}}q_{ij}y_{ij}^{s}$ (17a)
$\sum_{j\in\mathcal{N_{T}}}y_{jn}^{s}-\sum_{j\in\mathcal{N_{T}}}y_{nj}^{s}=b_{n}$ $\quad\forall\ n\in\mathcal{N_{C}}$ (17b)
$\sum_{j\in\mathcal{N_{T}}}y_{fj}^{s}-\sum_{j\in\mathcal{N_{T}}}y_{jf}^{s}\leq\xi_{x}^{fs}$ $\quad\forall\ f\in\mathcal{N_{F}}$ (17c)
$y_{ij}^{s}\leq Bz_{e(i,j)}$ $\quad\forall\ (i,j)\in\mathcal{A}\backslash\mathcal{A}_{m}$ (17d)
$\xi_{x}^{fs}=t(x,\vartheta^{s})$ $\quad\forall\ f\in\mathcal{N_{F}}$ (17e)
$y_{ij}^{s}\geq 0$ $\quad\forall\ (i,j)\in\mathcal{A}.$ (17f)
In the first stage, the objective function (16a) minimizes the expected value of the transformed second stage program. We also have Constraints (16b) and (16c), which specify the selection of one level of protection for each facility and ensure that the cost budget for investing in facility protection and edge installation, respectively, is respected. The objective function (17a) of the second stage minimizes the average cost of transporting goods from facilities to clients. In addition, the second stage contains Constraints (17b)–(17e). These ensure flow conservation at client nodes (Constraints (17b)), the use only of available capacity (Constraints (17c)), the existence of flow only in installed arcs (Constraints (17d)), and a suitable definition of the endogenous realizations $\xi_{x}^{fs}$ (Constraints (17e)). We show the transformed SAA Program (7) of the NDFPP in Appendix 11.
In the following sections, we describe the different variants of this problem according to the considered decision-dependent distribution of $\xi_{x}^{fs}$. In other words, we detail Constraints (17e) and define the $\vartheta^{s}$. For the discrete endogenous distributions, we assume the post-disruption capacity of facilities to have $|\mathcal{W}|$ possible values, where $\mathcal{W}=\{0,\ldots,W\}$ is an ordered index set such that level 00 represents no capacity and level $W$ represents maximum capacity. The parameter $\nu_{w}$, with $w\in\mathcal{W}$, is the capacity at level $w$ such that $\nu_{w+1}=\nu_{w}+\overline{\nu}$; that is, $\overline{\nu}$ is the increase in capacity per unit increase in capacity level.
5.1.1 Discrete Selection of Distributions
First, we consider the discrete selection of distributions introduced by Bhuiyan et al. (2020). If a protection level $p\in\mathcal{P}$ is installed at facility $f\in\mathcal{N_{F}}$, then its capacity level is considered to follow $Bin(W,\overline{\phi}_{f}^{dp})$ when event $d\in\mathcal{D}$ is realized. The parameter $\overline{\phi}_{f}^{dp}$ is the probability of successfully increasing the capacity level by one. We follow the technique in Section 4.1. For facility $f\in\mathcal{N_{F}}$, protection $p\in\mathcal{P}$, and scenario $s\in\mathcal{S}$, we define $\vartheta^{s}_{fp}$ as the value of the realization of the capacity level, which follows the $Bin(W,\overline{\phi}_{f}^{dp})$. In this case, no extra variables are needed. Constraints (17e) are
$\xi_{x}^{fs}=\sum_{p\in\mathcal{P}}\vartheta^{s}_{fp}x_{f}^{p}$ $\forall\ f\in\mathcal{N_{F}}.$ (18)
Finally, this variant of the considered problem is denoted NDFPP-Selection.
5.1.2 Binomial Distribution
We assume that the network-wide protection allocation affects the capacity level distribution of all facilities. Thus, we aim to represent situations in which facilities collaborate by sharing resources, situations in which facilities intercommunicate with transmission lines (e.g., power systems), or situations where failures propagate in the network (e.g., the spread of diseases). We consider that the capacity level of a facility $f\in\mathcal{N_{F}}$ follows $Bin(W,{\phi}_{f}^{d})$ when event $d\in\mathcal{D}$ occurs, where ${\phi}_{f}^{d}\in[0,1]$ is a variable computed as:
${\phi}_{f}^{d}=\frac{1}{\rho}\sum_{i\in\mathcal{N_{F}}}\sum_{p\in\mathcal{P}}\widetilde{\phi}_{fi}^{dp}x_{i}^{p}$ $\forall\ d\in\mathcal{D},\quad f\in\mathcal{N_{F}},$ (19)
where the parameter $\widetilde{\phi}_{fi}^{dp}$ is the impact on the capacity level at facility $f\in\mathcal{N_{F}}$ of installing protection level $p\in\mathcal{P}$ at facility $i\in\mathcal{N_{F}}$ when event $d$ occurs, and $\rho$ is a normalization parameter to guarantee that ${\phi}_{f}^{d}\in[0,1]$. Hence, we adopt convolution for the Binomial distribution in Section 4.4; that is, $\vartheta_{fw}^{s}$ is the realization of a random variable following $U(0,1)$ for scenario $s\in\mathcal{S}$, facility $f\in\mathcal{N_{F}}$, and capacity level $w\in\{1,\ldots,W\}$. The variables $\upsilon^{s}_{fw}\in\{0,1\}$ are equal to 1 if $\vartheta_{fw}^{s}<{\phi}^{\hat{d}(s)}_{f}$ and 0 otherwise for $f\in\mathcal{N_{F}}$ and $w\in\{1,\ldots,W\}$. Constraints (17e) are equivalent to
$\vartheta_{fw}^{s}\upsilon^{s}_{fw}\leq{\phi}_{f}^{\hat{d}(s)}-\epsilon$ $\forall\ f\in\mathcal{N_{F}},\quad w\in\{1,\ldots,W\}$ (20)
$\upsilon^{s}_{fw}\geq{\phi}_{f}^{\hat{d}(s)}-\vartheta_{fw}^{s}$ $\forall\ f\in\mathcal{N_{F}},\quad w\in\{1,\ldots,W\}$ (21)
$\xi_{x}^{fs}=\overline{\nu}\left(\sum_{w=1}^{W}\upsilon_{fw}^{s}\right)$ $\forall\ f\in\mathcal{N_{F}}.$ (22)
Constraints (20) and (21) guarantee the adequate definition of variables $\upsilon^{s}_{fw}$. Constraints (22) determine the capacity value of facilities in every scenario. This variant of the considered problem is denoted NDFPP-Binomial. Note that the original program of this variant is nonlinear nonconvex.
5.1.3 Discrete Distribution
We continue to assume that network-wide protection influences the capacity levels of all facilities. We consider that the random variables follow a discrete distribution inspired by the multinomial logit models in discrete choice theory (Bierlaire 1998). Therefore, the probability $\hat{p}_{df}^{w}(u)$ of a facility $f\in\mathcal{N_{F}}$ having capacity level $w\in\mathcal{W}$ after the occurrence of an event $d\in\mathcal{D}$ is defined as
$\hat{p}_{df}^{w}(u)=\frac{u_{df}^{w}}{\sum_{j\in\mathcal{W}}u_{df}^{j}}$ $\ \forall\ d\in\mathcal{D},\quad f\in\mathcal{N_{F}},\quad w\in\mathcal{W},$ (23)
where $u_{df}^{w}\geq 0$ are auxiliary variables expressing the likelihood of facility $f\in\mathcal{N_{F}}$ having capacity level $w\in W$ when event $d\in\mathcal{D}$ occurs. These variables are
$u_{df}^{w}=\sum_{i\in\mathcal{N_{F}}}\sum_{p\in\mathcal{P}}\widetilde{u}_{fi}^{dpw}x_{i}^{p}$ $\forall\ d\in\mathcal{D},\quad f\in\mathcal{N_{F}},\quad w\in\mathcal{W},$ (24)
where the parameter $\widetilde{u}_{fi}^{dpw}$ is the impact of installing protection $p\in\mathcal{P}$ at facility $i\in\mathcal{N}_{F}$ on the likelihood of capacity level $w\in\mathcal{W}$ at facility $f\in\mathcal{N_{F}}$ when event $d\in\mathcal{D}$ occurs. Thus, we adapt the inverse transformation in Section 4.2.1. For each facility $f\in\mathcal{N_{F}}$ and scenario $s\in\mathcal{S}$, the realization $\vartheta_{f}^{s}$ follows $U(0,1)$. The variables $\theta_{fw}^{s}\in\{0,1\}$ equal one if the capacity level $w\in\mathcal{W}$ is selected for facility $f\in\mathcal{N_{F}}$ in scenario $s\in\mathcal{S}$ and zero otherwise. We define the variables $\pi_{fw}^{s}\in\{0,1\}$, which are equal to $\sum_{j=0}^{w}\theta_{fj}^{s}$, with $\pi_{fW}^{s}$ set to one. This results in the following constraints:
$\pi_{fw}^{s}\sum_{w\in\mathcal{W}}u_{\hat{d}(s)f}^{w}\leq\frac{1}{\vartheta_{f}^{s}}\sum_{j=0}^{w}u_{\hat{d}(s)f}^{j}-\epsilon$ $\forall\ f\in\mathcal{N_{F}},\quad w\in\{0,\ldots,W-1\}$ (25)
$(1-\pi_{f,w-1}^{s})\sum_{w\in\mathcal{W}}u_{\hat{d}(s)f}^{w}\leq\frac{1}{(1-\vartheta_{f}^{s})}\sum_{j=w}^{W}u_{\hat{d}(s)f}^{j}$ $\forall\ f\in\mathcal{N_{F}},\quad w\in\{1,\ldots,W\}$ (26)
$\pi_{fw}^{s}=\pi_{f,w+1}^{s}-\theta_{f,w+1}^{s}$ $\forall\ f\in\mathcal{N_{F}},\quad w\in\{0,\ldots,W-1\}$ (27)
$\sum_{w\in\mathcal{W}}\theta_{fw}^{s}=1$ $\forall\ f\in\mathcal{N_{F}}$ (28)
$\xi_{x}^{fs}=\sum_{w\in\mathcal{W}}\nu_{w}\theta_{fw}^{s}$ $\forall\ f\in\mathcal{N_{F}}.$ (29)
Constraints (25)–(28) are, respectively, the adaptation of Constraints (11)–(14) for the NDFPP. They ensure the correct inverse transformation of $\vartheta_{f}^{s}$ according to the probability distribution defined by Equation (23). Constraints (29) delineate the capacity of facilities according to the capacity level obtained from the proposed transformation. Finally, Constraints (25) and (26) have bilinear terms involving binary variables. To satisfy Assumption 3, we apply the exact McCormick envelope (see Appendix 12 for the linear reformulation). This variant is denoted NDFPP-Discrete. Finally, the original program in this case is also nonlinear nonconvex.
5.1.4 Normal Distribution
We also assume that network-wide protection affects the facilities’ capacity level. We consider that the capacity level of facility $f\in\mathcal{N_{F}}$ follows $N(\mu_{f}^{d},{(\widetilde{\sigma}_{f}^{d})}^{2})$ when event $d\in\mathcal{D}$ occurs. The standard deviation $\widetilde{\sigma}_{f}^{d}\geq 0$ is a parameter and the mean $\mu_{f}^{d}\geq 0$ is a decision variable such that
$\mu_{f}^{d}=\frac{1}{\rho}\sum_{i\in\mathcal{N_{F}}}\sum_{p\in\mathcal{P}}\widetilde{\mu}_{fi}^{dp}x_{i}^{p}$ $\forall\ d\in\mathcal{D},\quad f\in\mathcal{N_{F}}.$ (30)
The parameter $\widetilde{\mu}_{fi}^{dp}$ is the impact of installing protection $p\in\mathcal{P}$ at facility $i\in\mathcal{N}_{F}$ on the mean of the capacity level of facility $f\in\mathcal{N}_{F}$ when event $d\in\mathcal{D}$ occurs, and $\rho$ is a normalization value. We assume that the capacity is always positive (i.e., $4\widetilde{\sigma}_{f}^{d}\leq\mu_{f}^{d}$ for all feasible decisions). Hence, we use the standardization method in Section 4.3, so the realization $\vartheta_{f}^{s}$ follows $N(0,1)$ for each facility $f\in\mathcal{N_{F}}$ and scenario $s\in\mathcal{S}$. Constraints (17e) are modeled as
$\xi^{fs}_{x}=\mu_{f}^{\hat{d}(s)}+\vartheta_{f}^{s}\sigma_{f}^{\hat{d}(s)}$ $\forall\ f\in\mathcal{N_{F}}.$ (31)
This problem variant is denoted NDFPP-Normal, and its original program is nonlinear nonconvex.
6 Experimental Results
This section validates the application of our methodology for addressing 2-SMIPD. First, Section 6.1 summarizes the instance generation. Section 6.2 then compares the transformed program for NDFPP-Selection solved with the SAA and the DEP of its original program, specifically, the MILP reformulation obtained by auxiliary linearizations in Bhuiyan et al. (2020). The DEP of the original program for the NDFPP-Binomial, NDFPP-Discrete, and NDFPP-Normal is nonlinear nonconvex (see Section 5). Thus, Section 6.3 only explores the performance of the estimators obtained by the SAA applied to our transformed program. Finally, we investigate the value of the stochastic solution (VSS) of the NDFPP in Section 6.4. Note that different methods can be used to solve our obtained two-stage program with exogenous uncertainty. We apply the SAA given its simplicity and desirable properties (see Section 3.1).
The instances of the SAA program, the DEP, the EV problem, and the expected result of the EV solution (EEV) for computing the VSS were coded in C++ and solved directly with Gurobi 10.0.3. All experiments were processed on the Beluga cluster of the Compute Canada network on an Intel Gold 6148 Skylake @ 2.4 GHz with a maximum of 100 GB. We limited Gurobi to using one thread. The details on the application of the SAA are given in Appendix 8. The code is available on GitHub (https://github.com/mariabazotte/saa_endog).
6.1 Instance Generation
We follow a procedure similar to that of Bhuiyan et al. (2020) for network generation. We use the latitude, longitude, and population data of cities in the southeastern United States from the basic US Cities dataset, which is available at https://simplemaps.com/data/us-cities. First, we select cities with a population exceeding a predefined threshold as the nodes of the network. Then, departing from Bhuiyan et al. (2020), we use Delaunay triangulation to define the edges of the network. Thus, we do not restrict the edge lengths and the number of neighbors for nodes. We investigate five networks, with $|\mathcal{N}_{\mathcal{T}}^{0}|=15$, 25, 30, 39, and 48. The thresholds and number of edges for these networks are presented in Appendix 15. The transportation cost $q_{ij}$ of arcs $(i,j)\in\mathcal{A}\backslash\mathcal{A}_{m}$ is the geometric distance $\varrho_{e(i,j)}$ between cities $i$ and $j$. The demands are set to the node population divided by $10^{4}$. We set the dummy facility cost multiplier $a$ to $10$. Following Bhuiyan et al. (2020), we solve the capacitated facility location and network design problem from Melkote and Daskin (2001) to define the set of facilities. For this, we use the median home value from www.zillow.com, 2023, and we set the node capacity to ${\sum_{n\in\mathcal{N}_{\mathcal{T}}^{0}}b_{n}}/{0.9|\mathcal{N}_{\mathcal{T}}^{0}|}$. We consider the values 4 and 5 for $|\mathcal{N}_{\mathcal{F}}|$. This code is available from the same GitHub project.
[…]
Analogously to Bhuiyan et al. (2020), we define the probability of success of increasing the capacity level by one $\overline{\phi}_{f}^{dp}$ as $({0.95\times c_{f}^{p}}/{c_{f}^{P}})^{{\hat{l}(f,d)}/{L}}$, where $\hat{l}(f,d)$ is the intensity level in $\mathcal{L}$ at which facility $f\in\mathcal{N_{F}}$ is impacted by disruption $d\in\mathcal{D}_{0}$. However, we consider a 5% risk of protection malfunction on the success of one capacity level. For the no-disruption ($nd$) event, we assign $\overline{\phi}_{f}^{nd,p}$ the value $\eta({0.95\times c_{f}^{p}}/{c_{f}^{P}})+(1-\eta)$, where $\eta=0.7$ represents the degree to which the protection decisions impact the uncertainty in the capacity level for this event. Bhuiyan et al. (2020) consider that facilities are not impacted by the no-disruption ($nd$) event and by disruption events occurring in low-intensity-level regions. Thus, we have instances with a larger number of scenarios.
For the NDFPP-Binomial and NDFPP-Discrete, we define $\widetilde{\phi}_{fi}^{dp}$ and $\widetilde{u}_{fi}^{dpw}$ as $\overline{\phi}_{i}^{dp}$ if $f=i$, and as $0.3\cdot\overline{\phi}_{i}^{dp}$ otherwise. For the NDFPP-Binomial and NDFPP-Normal, we define $\rho=1+0.3\cdot(|\mathcal{N_{F}}|-1)$. For the NDFPP-Normal, we set $\widetilde{\mu}_{fi}^{dp}$ to $\overline{\mu}_{f}^{dp}$ if $f=i$, and to $0.3\cdot\overline{\mu}_{f}^{dp}$ otherwise. The parameter $\overline{\mu}_{f}^{dp}$ is defined as $\nu_{\text{max}}\left[1-\frac{(P+1-p)[L+\hat{l}(f,d)+2]}{2(P+1)(L+1)}\right]$ for each protection level $p\in\mathcal{P}$, where $\hat{l}(f,d)$ returns zero for the no-disruption event ($nd$). Thus, the mean increases with increasing protection and decreases with increasing intensity levels. The standard deviation $\widetilde{\sigma}^{d}_{f}$ is given by $\frac{1}{\rho}\left[\overline{\sigma}_{f}^{d}+0.3\cdot\sum_{i\in\mathcal{N}_{F}\backslash\{f\}}\overline{\sigma}_{i}^{d}\right]$ where $\overline{\sigma}_{f}^{d}$ is defined as $\nu_{\text{max}}\left[\frac{L+\hat{l}(f,d)+2}{8(P+1)(L+1)}\right]$. Hence, it increases with increasing intensity levels and is independent of protection. This way, $0\leq 4\overline{\sigma}_{f}^{d}\leq\overline{\mu}_{f}^{d1}\leq\cdots\leq\overline{\mu}_{f}^{dP}$.
Given that the costs $c_{f}^{p}$ are randomly generated, we consider five different seeds per instance configuration, (network, number of facilities, and capacity level). Thus, for NDFPP-Selection, NDFPP-Binomial, and NDFPP-Discrete, we have $5\times 2\times 3=30$ instances per seed (150 in total). For NDFPP-Normal, we have $5\times 2=10$ instances per seed (50 in total).
6.2 Comparative results for NDFPP-Selection
The DEP of the original program is formulated in Bhuiyan et al. (2020) as an MILP obtained by linearizing the probability function (4). To solve this MILP, we implemented the same methodology as Bhuiyan et al. (2020): an L-shaped decomposition accelerated with nondominated cuts and valid inequalities. We compare this approach with solving our transformed program with the SAA. Note that we cannot apply the SAA to the original program as it has endogenous uncertainty.
[…]
0010101010202020203030303040404040505050506060606070707070808080809090909010010010010000101010102020202030303030404040405050505060606060707070708080808090909090100100100100Instances (%)SAA Transformed program $\left(\overline{gap}(\overline{x})(\%)\right)$DEP Original program $\left(GAP(\%)\right)$
Comparison of the optimality gap obtained by the SAA for the transformed program and the DEP of the original program.
Table 6.2 shows the experimental results of the 4- and 5-facility instances. In the table, $|K|$ and $|K_{t}|$ are the set of scenarios in the DEP of the original program (Program (2)) and in the DEP of our transformed Program (6), respectively. Moreover, $GAP(\%)$ is the optimality gap for the DEP. For the DEP, all runs reached the time limit of 2 hours. We followed the procedure explained in Section 3.1 for the SAA to obtain the feasible solution $\overline{x}$ and its estimators (see Appendix 8 for details). Even though the number of scenarios $|K_{t}|$ of the transformed program is extremely large (i.e.,$|\mathcal{D}|{|\mathcal{W}|}^{|\mathcal{N_{F}}||\mathcal{P}|}$), the number $|K|$ of scenarios of the DEP is small (i.e., $|\mathcal{D}|{|\mathcal{W}|}^{|\mathcal{N_{F}}|}$) for the considered instances. Thus, we can compute the real stochastic value of the feasible solution $\overline{x}$, i.e., $v(\overline{x})=c^{\top}\overline{x}+g(\overline{x})$. For the SAA, we solve $M=50$ problems with $N=750$ sample scenarios because these values sufficed to obtain estimators with low variance and feasible, high-quality solutions. We present the over-estimator of the relative optimality gap, i.e., $\overline{gap}(\overline{x})(\%)={100\%\times[v(\overline{x})-\overline{v}_{M}^{N}]}/{v(\overline{x})}$. Because the lower-bound estimator $\overline{v}_{M}^{N}$ is a random variable, this over-estimator of the optimality gap may assume negative values (Verweij et al. 2003). The computational time of the SAA is in seconds. For all runs, the SAA finished before the time limit of 2 hours. For comparison, we display the gap between the upper-bound values obtained by the proposed method and the DEP, i.e., $GAP_{UB}(\%)={100\%\times[UB-v(\overline{x})]}/{UB}$, and the gap between their lower bounds, i.e., $GAP_{LB}(\%)={100\%\times\left(LB-\overline{v}_{M}^{N}\right)}/{LB}$, where $UB$ and $LB$ are the upper and lower bounds for the DEP, respectively. We summarize our main findings in Results 1 and 2.
Result 1
The proposed method produces a feasible solution $\overline{x}$ with a considerably low over-estimator of the optimality gap, $\overline{gap}(\overline{x})(\%)$ is between $0\%$ and $1\%$ for all instances. The lower bound $\overline{v}_{M}^{N}$ and the optimality gap estimators have low variability, as shown by the standard deviation $\hat{\sigma}_{N,M}$ of the lower-bound estimator, which is inferior to $0.5\%$ of $\overline{v}_{M}^{N}$ for all instances (see Appendix 16 for details). For the considered instances, the computational time required by the proposed method was inferior to 2000s in 87% of the cases, with an average of around 990s.
Result 2
The proposed method’s over-estimator of the optimality gap is significantly smaller than the DEP’s optimality gap (the difference is around $24\%$ on average). The proposed method obtained these optimality gaps in significantly less time than the DEP, and, within the time limit, consistently found feasible solutions with slightly better values than those from the DEP.
Result 1 highlights the good performance of the proposed method. Result 2 shows that it outperformed the DEP. Figure 6.2, which shows the optimality gap for the DEP and the proposed approach, corroborates this analysis. The proposed method obtained optimality gaps close to $0\%$ for all instances, whereas the DEP obtained optimality gaps inferior to $10\%$ only for $22\%$ of the instances. The DEP obtained optimality gaps of $10\%$$20\%$ for around $56\%$ of the instances, of $20\%$$30\%$ for around $8\%$ of the instances, and of $100\%$ for around $14\%$ of the instances.
As previously stated, the transformed Program (6) has more scenarios than the DEP of the original program. Nonetheless, we apply the SAA to our transformed program to efficiently solve different problems with fewer scenarios. The proposed approach obtained high-quality solutions by sampling only $N=750$ scenarios for each of the $M=50$ SAA problems, which, in most cases, is less than the number $|K|$ of scenarios of the DEP. Despite the number of scenarios, the proposed method solves mathematical programs smaller than those obtained with the DEP (in terms of their number of variables and constraints). The DEP’s MILP formulation contains numerous big-M constraints due to the linearization of the probability function (4), which leads to weak linear relaxations. In contrast, the transformation function applied to obtain the NDFPP-Selection formulation (see Sections 4.1 and 5.1.1) includes no extra variables or constraints.
6.3 Analysis for NDFPP-Binomial, NDFPP-Discrete, and NDFPP-Normal
In this section, we analyze the statistical estimators and feasible solution $\overline{x}$ of the SAA applied to our transformed program for NDFPP-Binomial, NDFPP-Discrete, and NDFPP-Normal. In these cases, the DEP of the original program (Program (2) for discrete distributions) is nonlinear nonconvex, and the linearization of Bhuiyan et al. (2020) cannot be applied. Moreover, note that our transformed Program (6) for these variants has continuous random variables, making the explicit enumeration of scenarios impossible and requiring the implementation of sampling techniques.
The DEP of the original program has a small number $|K|$ of scenarios for the NDFPP-Binomial and NDFPP-Discrete instances. Thus, we compute $v(\overline{x})$ for these cases. The number of scenarios of the DEP of the original program is infinite for NDFPP-Normal because the endogenous random variables are continuous. Consequently, we compute the estimator $\hat{v}_{N{{}^{\prime}}}(\overline{x})$ of $v(\overline{x})$. Table 6.3 lists the experimental results of the 4- and 5-facility instances. The over-estimator of the optimality gap $\overline{gap}(\overline{x})(\%)$ is defined as in the previous section for NDFPP-Binomial and NDFPP-Discrete. For NDFPP-Normal, $\overline{gap}(\overline{x})(\%)={100\%\times[\hat{v}_{N{{}^{\prime}}}(\overline{x})-\overline{v}_{M}^{N}]}/{\hat{v}_{N{{}^{\prime}}}(\overline{x})}$. We set $M=50$ and $N=750$ for the same reason as in the previous section. For NDFPP-Normal, we use $N{{}^{\prime}}=5\times 10^{4}$. The SAA finished for all instances before reaching 3 hours. The result is as follows:
Result 3
The proposed method obtained a feasible solution $\overline{x}$ with a significantly low statistical over-estimator of the optimality gap $\overline{gap}(\overline{x})(\%)$ between $-1\%$ and $1\%$ for all instances of the problem versions. The SAA’s estimators have high confidence, as shown by the standard deviation $\hat{\sigma}_{N,M}$ of the lower-bound estimator, which is inferior to $0.5\%$ of $\overline{v}_{M}^{N}$ for all cases, and the standard deviation $\hat{\sigma}_{N{{}^{\prime}}}(\overline{x})$ of the upper bound estimator, which is inferior to $0.2\%$ of $\hat{v}_{N{{}^{\prime}}}(\overline{x})$ for all NDFPP-Normal instances (see Appendix 17 for details). The computational time was, on average, around 1500s for NDFPP-Binomial, 3000s for NDFPP-Discrete, and 1200s for NDFPP-Normal.
[…]
The proposed method found high-quality solutions and estimators for the considered problem variants. The computational time of this approach for the NDFPP-Binomial, NDFPP-Discrete, and NDFPP-Normal variants exceeded that for the NDFPP-Selection variant. Solving the NDFPP-Discrete variant required the longest computational time. The transformation function used in this case (i.e., inverse transformation for discrete random variables; see Section 4.2.1) includes numerous binary variables and big-M constraints in the formulation. The initial formulation of NDFPP-Discrete contained bilinear terms, which were linearized by adding more big-M constraints in the formulation. For the same reason, the computational time of the proposed method for NDFPP-Binomial exceeded that for NDFPP-Selection. The computational time of the proposed method for solving NDFPP-Discrete increases with the number of capacity levels (possible realizations of the random variables). Therefore, a trade-off exists between how sample size affects the performance and confidence of the SAA’s statistical estimators and the associated computational time. For NDFPP-Normal, the number of evaluation scenarios $N{{}^{\prime}}$ is considerably large, which increases the computational time of the proposed method. Moreover, the standard deviation of the optimality gap is slightly larger than for the remaining versions because it includes the variability of $\hat{v}_{N{{}^{\prime}}}(\overline{x})$.
6.4 Value of the Stochastic Solution
This section analyzes the relative VSS of all problem variants. The relative VSS is defined as ${(\textrm{EEV}-v^{*})}/{\textrm{EEV}}$, where $v^{*}$ is the optimal value of the recourse problem, and the EEV is the expected result of using the EV solution. Thus, $\textrm{EEV}=v(\hat{x}^{*})=c^{\top}\hat{x}^{*}+g(\hat{x}^{*})$, where $\hat{x}^{*}$ is the optimal solution of the EV. In Appendix 13, we detail the EVs (i.e., the deterministic problems considering the average case scenario) according to the problem variants. We use the feasible solution $\overline{x}$ obtained by the SAA applied to our transformed program to compute under estimators of the VSS. We have the under estimator $\textrm{VSS}_{1}={[\textrm{EEV}-v(\overline{x})]}/{\textrm{EEV}}$ for NDFPP-Selection, NDFPP-Binomial, and NDFPP-Discrete, for which we can enumerate the DEP scenarios. For NDFPP-Normal, we compute $\textrm{VSS}_{2}={[\overline{\textrm{EEV}}-\hat{v}_{N{{}^{\prime}}}(\overline{x})]}/{\overline{\textrm{EEV}}}$, where $\overline{\textrm{EEV}}=\hat{v}_{N{{}^{\prime}}}(\hat{x}^{*})$ and $N{{}^{\prime}}=5\times 10^{4}$. Table 6.4 summarizes the under estimators of the VSS for the 4- and 5-facility instances.
[…]
The proposed method obtained better solutions than the EV in all cases. Thus, solving the recourse problem instead of the EV is also advantageous in the context of endogenous uncertainty, where the distribution defined by $\hat{x}^{*}$ is not the same as that defined by $\overline{x}$ or $x^{*}$. Different instances of NDFPP-Selection showed a high VSS, around $30\%$ or more. Some instances of NDFPP-Binomial and NDFPP-Discrete also presented high VSSs, whereas others had VSSs of $1\%$$4\%$. Moreover, although the VSS for NDFPP-Normal approached $0\%$ for the 5-facility instances, it hovered around $3\%$$6\%$ for the 4-facility instances. Although the VSS is primarily instance-dependent, it is also affected by the problem (e.g., NDFPP-Selection and NDFPP-Normal for 15-node instances).
7 Conclusions
We proposed a general method for addressing 2-SMIPD that uses the random variable transformation. Our methodology (i) works for a general class of programs with endogenous uncertainty; (ii) obtains a stochastic program with exogenous uncertainty for which solution procedures are well-studied; (iii) yields MILPs or MICPs with exogenous uncertainty in many cases; (iv) works for both discrete and continuous endogenous random variables; and (v) is applicable to problems with independent and correlated random variables. We investigated techniques to define appropriate transformations of random variables to treat cases where the decision-dependent random variables follow classical probability distributions. We showed the existence of a simple transformation for the case of the discrete selection of distributions, which is the most commonly studied in previous research. We also demonstrated the proposed methodology by applying it to the NDFPP, a broadly studied problem in the literature. In the computational experiments, the SAA applied to our transformed program produced feasible, high-quality solutions and high-confidence estimators in a short computational time.
Numerous research perspectives stem from this work. Future work could focus on (i) exploring additional techniques to define transformation functions; (ii) investigating valid inequalities for the discrete random variables inverse transformation; (iii) analyzing how endogenous uncertainty affects the definition of the VSS; (iv) examining stochastic programs with endogenous random variables for which the probability distribution is not known a priori; and (v) extending the NDFPP to account for assumptions commonly investigated in previous research, such as post-disruption edge failure, protection (Viswanath et al. 2004, Flach 2010, Peeta et al. 2010, Prestwich et al. 2014, Du and Peeta 2014), and facility location (Li et al. 2021, 2022).
[…]
This research was funded by the SCALE-AI Chair in Data Driven Supply Chains, the FRQ-IVADO Research Chair in Data Science for Combinatorial Game Theory, and NSERC Grant No. 2024-04051. It was also enabled by support provided by Calcul Québec (www.calculquebec.ca) and the Digital Research Alliance of Canada (https://alliancecan.ca/).