### v4

# Maria Bazotte, Margarida Carvalho, and Thibaut Vidal

[…]
Maria Bazotte, Margarida Carvalho, and Thibaut Vidal
[…]
Solving Programs with Endogenous Uncertainty via Random Variable Transformation
[…]
Solving Two-Stage Programs with Endogenous Uncertainty via Random Variable Transformation
[…]
Real-world decision-making problems often involve decision-dependent uncertainty, where the probability distribution of the random vector depends on the model’s decisions. Few studies focus on two-stage stochastic programs with this type of endogenous uncertainty, and those that do lack general methodologies. We propose a general method for solving a class of these programs based on random variable transformation, a technique widely employed in probability and statistics. The random variable transformation converts a stochastic program with endogenous uncertainty (original program) into an equivalent stochastic program with decision-independent uncertainty (transformed program), for which solution procedures are well-studied. Additionally, endogenous uncertainty usually leads to nonlinear nonconvex programs, which are theoretically intractable. Nonetheless, we show that, for some classical endogenous distributions, the proposed method yields mixed-integer linear or convex programs with exogenous uncertainty. We validate this method by applying it to a network design and facility-protection problem, considering distinct decision-dependent distributions for the random variables. While the original formulation of this problem is nonlinear nonconvex for most endogenous distributions, the proposed method transforms it into mixed-integer linear programs with exogenous uncertainty. We solve these transformed programs with the sample average approximation method. We highlight the superior performance of our approach compared to solving the original program in the case a mixed-integer linear formulation of this program exists.
[…]
decision-dependent or endogenous uncertainty, network design and facility protection, random variable transformation, two-stage stochastic programs
1 Introduction
Uncertainty is an inherent characteristic of most real-world decision-making problems (Küçükyavuz and Sen 2017). Stochastic programming is a well-known framework in mathematical programming to model optimization problems with uncertain parameters; see, e.g., (Shapiro et al. 2021) for a summary of the theoretical aspects and the most popular algorithms. Research in this area has focused mainly on exogenous or decision-independent uncertainty (Li and Grossmann 2021), where the random vector is independent of the model decisions. In contrast, endogenous or decision-dependent uncertainty, where the random vector depends on the decisions, has received less attention, although it is prevalent in critical applications such as pre-disaster supply chain management (Peeta et al. 2010, Bhuiyan et al. 2020), and production planning (Ekin 2018).
In the context of endogenous uncertainty, different types of decision dependency may occur (Jonsbrten et al. 1998). Goel and Grossmann (2006) propose classifying decision dependency into two types: For problems with Type 1 endogenous uncertainty or decision-dependent probability, decisions impact the underlying probability distribution of the random parameters. Although little research focuses on this type of uncertainty, as noted by Li and Grossmann (2021), it deals with various problems of practical relevance, such as the design of communication networks, power plants, and warehouses and resource allocation in queuing systems (Pflug 1990). With Type 2 endogenous uncertainty, decisions influence the time at which the realization of the random variables is totally or partially revealed. Thus, with Type 2 endogenous uncertainty, the scenario tree depends on the decisions (Jonsbrten et al. 1998). Problems with Type 2 uncertainty have received more attention than those with Type 1, being mostly multistage programs. Li and Grossmann (2021) provide a concise review of methodologies and algorithms for multistage programs with Type 2 endogenous uncertainty. In the present work, we focus on Type 1 endogenous uncertainty, in particular, on two-stage stochastic mixed-integer linear programs with decision-dependent probability (2-SMIPD), in which the probability distribution depends on the first-stage decisions. In what follows, we refer to Type 1 endogenous uncertainty as endogenous uncertainty.
The mathematical properties of a stochastic program, such as convexity, are strongly linked to whether its uncertainty is modeled as endogenous or exogenous. Consequently, solution procedures typically employed for stochastic programs with exogenous uncertainty become complex when dealing with endogenous uncertainty, necessitating a priori a comprehensive analysis of its underlying properties (Varaiya and Wets 1989). In fact, endogenous uncertainty usually leads to nonlinear nonconvex programs (Dupacová 2006), which are theoretically intractable (NP-hard). Moreover, few studies focus on general-purpose methodologies to solve programs with endogenous uncertainty sharing common properties (i.e., methodologies are typically problem-specific).
Contributions.
The main contribution of this work is a general method based on random variable transformation to address 2-SMIPD with discrete or continuous endogenous random variables. This method resolves two main issues in the literature: the lack of generalist methods and the difficulty of solving the usual nonconvex nonlinear programs with endogenous uncertainty. More specifically, the contributions of this work are as follows: (i) We define a transformation function to model the endogenous uncertainty by combining first-stage decisions with decision-independent random variables, thereby converting stochastic programs with endogenous uncertainty (i.e., original programs) into stochastic programs with exogenous uncertainty (i.e., transformed programs). (ii) We identify different techniques to define appropriate transformation functions for endogenous random variables following classical decision-dependent probability distributions, including a commonly studied case known as the “discrete selection of distributions”. (iii) We show that, for these distributions, the obtained transformed programs have a convex continuous relaxation. (iv) We investigate the implications of endogenous uncertainty when applying sampling methods, particularly the sample average approximation (SAA) method, to solve the transformed program. (v) We apply the proposed approach to the network design and facility protection problem (NDFPP), considering different decision-dependent distributions. (vi) Finally, for all endogenous distributions, we analyze the statistical estimators of the SAA applied to the transformed program (SAA-TP) of the NDFPP. In less than one hour on average, the SAA-TP finds feasible solutions with significantly small over-estimated optimality gaps (between $-1\%$ and $1\%$) and a high confidence interval for all decision-dependent distributions. For the NDFPP with most endogenous distributions, the deterministic equivalent of the original program (DE-OP) is nonlinear nonconvex. The exception is the discrete selection of distributions case, for which the DE-OP is a mixed-integer linear program (MILP) obtained by auxiliary linearizations (Bhuiyan et al. 2020). In a considerably shorter time, the SAA-TP obtains a better over-estimator of the optimality gap than the optimality gap obtained by solving this mixed-integer DE-OP (a difference of around $24\%$ on average). Solving the SAA-TP also obtains better solutions than when using an average scenario, i.e., the expected value (EV) problem.
Organization.
The remainder of this work is structured as follows: Section 2 formalizes the class 2-SMIPD and discusses related works on endogenous uncertainty. Section 3 presents our general methodology to obtain the transformed program and discusses the application of methods such as the SAA to solve this resulting program. Section 4 catalogs techniques to define transformations to various endogenous distributions, and Section 5 explains how to apply this approach to the NDFPP. Finally, Section 5.2 reports the experimental results, comparing the proposed method with other solution procedures, and Section 6 concludes our analyses.
2 Problem Definition and Related Works
We focus on two-stage stochastic mixed-integer linear programs with decision-dependent probability (2-SMIPD). Let $x\in\mathbb{R}^{n_{1}}_{+}$ be the vector of first-stage variables (i.e., the “here-and-now” decisions) and $\xi(\omega)\in\mathbb{R}^{U}$ a random vector that depends on the random events $\omega\in\Omega$. For the sake of simplicity, we also denote $\bm{\xi}=\xi(\omega)$ and designate $\xi$ as a realization of this random vector. We use $\mathbb{E}_{\bm{\xi}\mid x}[\cdot]$ to denote the expected value operator with respect to $\bm{\xi}$ given the first-stage variables $x$. Finally, $y(\omega)\in\mathbb{R}^{n_{2}}_{+}$ is the vector of second-stage variables given a realization of the random events. The standard mathematical formulation for the 2-SMIPD is:
$\min$ $c^{\top}x+\mathbb{E}_{\bm{\xi}\mid x}\left[Q\bm{(}x,\xi(\omega)\bm{)}\right]$ (1a)
s.t. $Ax\geq b\ ,\ x\in\mathcal{X}$ (1b)
where $\textrm{where }Q\bm{(}x,\xi(\omega)\bm{)}=\min\{q(\omega)^{\top}y(\omega):T(\omega)x+W(\omega)y(\omega)\geq h(\omega),y(\omega)\in\mathcal{Y}\}.$ (1c)
In Program (1), $\mathcal{X}\subseteq\mathbb{R}^{n_{1}}_{+}$ and $\mathcal{Y}\subseteq\mathbb{R}^{n_{2}}_{+}$ are two compact sets imposing non-negativity and continuous, integer, or binary restrictions. The first stage has the matrices $A$, $b$, and $c$ with appropriate dimensions. The second-stage Program (1c) is established after the realization of the random events $\omega\in\Omega$ and has the matrices $q(\omega)$, $h(\omega)$, $T(\omega)$, and $W(\omega)$ with possible random elements and appropriate dimensions. Thus, the random vector $\xi^{T}(\omega)$ combines these random elements, i.e., $\xi^{T}(\omega)=\bm{(}q(\omega)^{\top},h(\omega)^{\top},T(\omega),W(\omega)\bm{)}$, and its support (according to its unconditional distribution on $x$) is denoted by $\Xi\subseteq\mathbb{R}^{U}$. For the sake of clarity, we define the random vectors ${\bm{\xi}_{x}^{\top}=\bm{(}q_{x}(\omega)^{\top},h_{x}(\omega)^{\top},T_{x}(\omega),W_{x}(\omega)\bm{)}}$ for $x\in\mathcal{X}$, which follow the conditional probability distribution of the random vector $\bm{\xi}$ given $x$ and have support $\Xi_{x}\subseteq\Xi$. We also set ${g(x)=\mathbb{E}_{\bm{\xi}\mid x}[Q\bm{(}x,\xi(\omega)\bm{)}]}$.
The 2-SMIPD assumes that the probability distribution of $\bm{\xi}$ depends on the first-stage decisions $x\in\mathcal{X}$ (i.e., it computes the conditional expectation $\mathbb{E}_{\bm{\xi}|x}[\cdot]$ instead of the standard expectation $\mathbb{E}_{\bm{\xi}}[\cdot]$ used in the exogenous case). This endogenous probability influences the properties of the resulting formulation. For example, unlike the exogenous case, the convexity of the continuous relaxation of $Q\bm{(}x,\xi(\omega)\bm{)}$ in Program (1) does not necessarily yield the convexity of the continuous relaxation of $\mathbb{E}_{\bm{\xi}|x}[Q\bm{(}x,\xi(\omega)\bm{)}]$ (Varaiya and Wets 1989, Dupacová 2006). As a result, algorithms commonly used for solving two-stage programs with exogenous uncertainty may not be suitable for 2-SMIPD.
This work assumes that the conditional probability distribution is known a priori and defined by a distinct probability mass or density function for $x\in\mathcal{X}$ (i.e., $x$ is a parameter of the distribution). We justify this assumption based on its broad application in the literature, as reviewed in Section 2.2. We consider it a reasonable extension of the usual stochastic optimization assumption for programs with exogenous uncertainty (i.e., random vectors with known probability distributions).
For the sake of simplicity, we next discuss our reasoning assuming that the random vector $\bm{\xi}$ is discrete and has finite support $\Xi$; the process is analogous for the continuous case. We write the DE-OP in its extensive form (i.e., the deterministic equivalent of Program (1)):
[…]
where $K$ is the set of realizations of $\bm{\xi}$ (i.e., the set of scenarios defined by the support $\Xi$). For each scenario $k\in K$, $q_{k}$, $h_{k}$, $T_{k}$, and $W_{k}$ have appropriate dimensions, and $\xi_{k}^{\top}=(q_{k}^{\top},h_{k}^{\top},T_{k},W_{k})$. In addition, $p_{k}(x):\mathcal{X}\rightarrow[0,1]$ is a function returning the conditional probability of scenario $k\in K$ given the first-stage variables $x$, in other words, $\bm{\xi}_{x}$ follows the probability distribution defined by $p_{k}(x)$ for every $k\in K$. We remark that scenarios in $\Xi\backslash\Xi_{x}$ have zero probability according to $p_{k}(x)$.
2.1 Background of Decision-Dependent Probability Distribution
In Program (2), the probability function $p_{k}(x)$ multiplies the second-stage objective term $q_{k}^{\top}y_{k}$, leading to nonlinear terms. Handling this nonlinearity requires describing $p_{k}(x)$, which depends on the assumptions of $\bm{\xi}$. When the $U$ elements of the random vector $\bm{\xi}$ are independent, we have
$p_{k}(x)=\prod_{u=1}^{U}\hat{p}_{\hat{r}(u,k)}^{u}(x)\quad\forall\ k\in K,$ (3)
where $\hat{r}(u,k)$ returns the realization in $\mathcal{R}_{u}=\{1,\ldots,R_{u}\}$ of a random element $u\in\mathcal{U}=\{1,\ldots,U\}$ in scenario $k\in K$. In addition, $\hat{p}_{\hat{r}(u,k)}^{u}(x):\mathcal{X}\rightarrow[0,1]$ is a function on $x$ returning the conditional probability of the realization $\hat{r}(u,k)$ of the random element $u\in\mathcal{U}$ in scenario $k\in K$. The function $p_{k}(x)$ can assume other configurations that do not necessarily use $\hat{p}_{\hat{r}(u,k)}^{u}(x)$ but instead return the conditional probability of the whole random vector $\bm{\xi}$.
A commonly studied case where decisions involve selecting a probability distribution from a predefined set is named “discrete selection of distributions” (Dupacová 2006, Hellemo et al. 2018). In this case, the decision-dependent probability is represented by a probability distribution for each independent random element $u\in\mathcal{U}$ from a set $\mathcal{D}_{u}=\{1,\ldots,D_{u}\}$ of available distributions, which can be either discrete or continuous. For discrete endogenous random variables, we have ${\hat{p}_{\hat{r}(u,k)}^{u}(x)=\sum_{d\in\mathcal{D}_{u}}\overline{p}_{\hat{r}(u,k)}^{ud}x_{u}^{d}}$ for every element $u\in\mathcal{U}$. Here, $x_{u}^{d}\in\{0,1\}$ is a subvector of $x\in\mathcal{X}$ representing the chosen distribution, such that $\sum_{d\in\mathcal{D}_{u}}x_{u}^{d}=1$ for every $u\in\mathcal{U}$. Moreover, the parameter $\overline{p}_{\hat{r}(u,k)}^{ud}$ represents the probability of realization $\hat{r}(u,k)$ of random component $u$ when distribution $d$ is selected. If a distribution from the set $\mathcal{D}=\{1,\ldots,D\}$ is chosen for the entire random vector $\bm{\xi}$ (i.e., for scenario $k\in K$), then $p_{k}(x)=\sum_{d\in\mathcal{D}}\overline{p}^{d}_{k}x^{d}$, where $\overline{p}^{d}_{k}$ is the probability of scenario $k$ under distribution $d$. Similarly, $x^{d}\in\{0,1\}$ are first-stage variables representing the selected distribution, such that $\sum_{d\in\mathcal{D}}x^{d}=1$. Additionally, Dupacová (2006) also categorizes problems involving the selection of parameters for a specific probability distribution as parameter selection. Hellemo et al. (2018) also introduce the probability-distortion case, where the chosen probability given $x$ is derived by combining a set of distributions. A general probability mass function $p_{k}(x)$ cannot be defined for these cases as it relies on additional assumptions.
2.2 Literature Review
Consistent with the assumptions of this work, the literature on decision-dependent probability focuses essentially on problems where the probability mass or density functions are known a priori. In this section, we review only works that use this assumption and that consider two-stage programs. We split our review into works on discrete and continuous random variables.
Discrete random variables.
For discrete random vectors, the elements are usually assumed to be independent. For the discrete selection of distributions, distinct applications of two-stage programs have been studied, such as network design protection (Viswanath et al. 2004, Flach 2010, Peeta et al. 2010, Prestwich et al. 2014), facility protection (Medal et al. 2016), facility location-protection (Li et al. 2021), fire risk management (Bhuiyan et al. 2019), supplier risk management (Zhou et al. 2022), facility protection and network design (Bhuiyan et al. 2020), and airline flight network expansion problems (Şafak et al. 2022). All these works used decision-dependent probabilities as in Equation (3), except for Şafak et al. (2022), who chose the probabilities of the entire random vector. These applications usually lead to large-scale instances with numerous scenarios.
Prestwich et al. (2014) and Bhuiyan et al. (2019) respectively obtained exact MILP reformulations of the DE-OP by using a linear scaling procedure and the probability chain technique described by O’Hanley et al. (2013). To solve large-scale problems, Prestwich et al. (2014) proposed a bundle method, and Medal et al. (2016), Bhuiyan et al. (2020), and Zhou et al. (2022) implemented the L-shaped decomposition for solving reformulations resulting from the probability chain technique. Usually, $T_{k}$ is a zero matrix in these two-stage programs (i.e., the second stage can be solved a priori). The exception is Zhou et al. (2022), who used an exact McCormick envelope (McCormick 1976) to obtain an MILP. These reformulations apply specifically to the case of the discrete selection of distributions and rely on the binary variables in $\hat{p}_{\hat{r}(u,k)}^{u}(x)$. However, they require numerous variables and big-M constraints. Thus, the practical use of these reformulations is limited to problems with a sufficiently small number of scenarios. Flach (2010) proposed a different approximate MILP reformulation for the DE-OP by using the binary variables in $\hat{p}_{\hat{r}(u,k)}^{u}(x)$ to reformulate $p_{k}(x)$ as a set of exponential and linear constraints. The authors ensured a predefined precision by approximating the exponential constraints by piecewise linear cuts included in a branch-and-cut scheme.
Various decision-dependent probability mass functions have also been studied, such as the binomial distribution (Karaesmen and Van Ryzin 2004, Green et al. 2013), linear interpolation (Du and Peeta 2014), a linear scaling and convex combination of a set of distributions (Hellemo et al. 2018), and the exponential function (Bhuiyan 2021). In most studies, the endogenous uncertainty in general cases of discrete random variables resulted in DE-OPs, which are mixed-integer nonlinear nonconvex programs (MINLPs) for which proving global optimality is harder than for MILPs or mixed-integer convex programs (MICPs). Even if $p_{k}(x)$ are convex or concave (see, e.g., Hellemo et al. 2018) or $\hat{p}_{\hat{r}(u,k)}^{u}(x)$ are convex or concave (see, e.g., Du and Peeta 2014, Bhuiyan 2021, Holzmann and Smith 2021, Homem-de Mello et al. 2022), the term $p_{k}(x)q_{k}^{\top}y_{k}$ could lead to an MINLP. Within this context, Krasko and Rebennack (2017) studied a two-stage preparedness hazard problem for debris-flow-hazard mitigation and solved the DE-OP with the BARON solver. Hellemo et al. (2018) solved the DE-OP of an energy investment problem using a generalized Benders decomposition (GBD) and the BARON solver. However, these approaches are limited to solving instances with few scenarios because the DE-OP in extensive form explicitly enumerates all scenarios.
Approximative approaches have been proposed for facility protection (Medal et al. 2016), sensors for holding water quality (Karaesmen and Van Ryzin 2004), and overbooking problems (Bhuiyan 2021). The objective function in these problems is submodular, allowing the design of greedy algorithms. However, approximation guarantees are generally low (e.g., around $63\%$ for Medal et al. (2016)). Li et al. (2021) and Li et al. (2022) used genetic algorithms to solve two-stage programs of a facility location-protection problem with a bilevel program as the first stage.
Viswanath et al. (2004), Peeta et al. (2010), and Du and Peeta (2014) used the SAA within iterative heuristic procedures to evaluate current first-stage solutions for two-stage network protection problems. Unfortunately, these procedures do not provide performance guarantees. The SAA was also used to estimate the performance of current solutions within stochastic gradient methods for overbooking (Karaesmen and Van Ryzin 2004) and appointment-scheduling problems (Homem-de Mello et al. 2022)Homem-de Mello et al. (2022) proved the local convergence of the algorithm, which is compatible with its nonconvexity. Finally, by considering a specific gradient estimator, Karaesmen and Van Ryzin (2004) proved the local convergence of the algorithm.
Flach (2010) used importance sampling to apply the SAA. The author sampled from a biased probability distribution and adjusted the weight of the objective function according to $p_{k}(x)$ and the biased distribution. However, importance sampling depends on the choice of the biased distribution (Shapiro et al. 2021, Chap. 5), and finding a unique biased distribution appropriate for all conditional distributions on $x$ could be difficult. Holzmann and Smith (2021) proposed an adapted SAA for bilevel shortest-path problems with randomized strategies. They sampled realizations from the continuous uniform distribution $U(0,1)$ and included big-M constraints in the model to transform these samples into realizations of the decision-dependent discrete distribution.
Continuous random variables.
To the best of our knowledge, only Ernst and Powell (1995), Ekin et al. (2017), Ekin (2018), and Hellemo et al. (2018) investigated continuous endogenous random variables. Ernst and Powell (1995) proposed a search procedure for solving a newsvendor problem with uncertain demand following a normal distribution with mean and variance selected by first-stage decisions. The augmented probability simulation (APS), in which decisions are also simulated, was used by Ekin et al. (2017) and Ekin (2018) for solving, respectively, a newsvendor problem with uncertain demand depending on stock level decisions and a production-planning problem with random yield depending on maintenance and production decisions. In both works, the dependency was modeled with a normal distribution with mean and variance defined according to the first-stage decisions. Finally, Hellemo et al. (2018) used the discretized function of the Kumaraswamy distribution and approximated the discretized function of the normal distribution. The discretization trades an increase in the number of scenarios for a confident discrete approximation. Approximations of the discretized function may also be needed when the cumulative distribution has no closed form or a complex closed form. In some applications, the search procedure cannot analyze the entire decision space, implying the need for a heuristic procedure. Finally, the APS remains limited to problems where the second stage is represented by analytical expressions.
3 Methodology
Our methodology is based on using random variable transformation, which is a classical topic in probability and statistics (see Hogg et al. 1977, Chap. 5), to model endogenous uncertainty. This section outlines Assumption 3, which is necessary for applying our methodology. Section 4 shows that this assumption is not restrictive by evidencing many cases in which it holds. {assumption} There exists (1) a random vector $\vartheta^{\top}(\omega)=\bm{(}\widetilde{q}(\omega)^{\top},\widetilde{h}(\omega)^{\top},\widetilde{T}(\omega),\widetilde{W}(\omega)\bm{)}$ with a known probability distribution that is independent of the first-stage decisions $x\in\mathcal{X}$ and (2) a transformation function $t(x,\bm{\vartheta})$, such that $\bm{\xi}_{x}=t\left(x,\bm{\vartheta}\right)$. Assumption 3 allows us to mimic the endogenous uncertainty by using a function $t(\cdot,\cdot)$ that combines the first-stage variables $x$ and the decision-independent random vector $\bm{\vartheta}$. For example, $t(\cdot,\cdot)$ can be the inverse transformation of $\bm{\xi}_{x}$ (Robert and Casella 1999, Chap. 2). Here, the random vector $\bm{\vartheta}$ does not necessarily have the same dimension as $\bm{\xi}$ or $\bm{\xi}_{x}$. Furthermore, the random elements of $\bm{\xi}_{x}$ may be dependent. For the sake of simplicity, we consider that $t(x,\bm{\vartheta})=(t_{1}\bm{(}x,\widetilde{q}(\omega)^{\top}\bm{)},t_{2}\bm{(}x,\widetilde{h}(\omega)\bm{)},t_{3}\bm{(}x,\widetilde{T}(\omega)\bm{)},t_{4}\bm{(}x,\widetilde{W}(\omega)\bm{)})$, where $t_{1}\bm{(}x,\widetilde{q}\left(\omega\right)^{\top}\bm{)}$, $t_{2}\bm{(}x,\widetilde{h}\left(\omega\right)\bm{)}$, $t_{3}\bm{(}x,\widetilde{T}\left(\omega\right)\bm{)}$, and $t_{4}\bm{(}x,\widetilde{W}\left(\omega\right)\bm{)}$ are transformation functions applied to the corresponding subvectors and matrices of $\bm{\vartheta}$. These functions return, respectively, the subvectors and matrices of $\bm{\xi}_{x}$, i.e., $q_{x}(\omega)^{\top}$, $h_{x}(\omega)$, $T_{x}(\omega)$, and $W_{x}(\omega)$. From Assumption 3, $t(.)$ and thus $t_{1}(.)$, $t_{2}(.)$, $t_{3}(.)$, and $t_{4}(.)$ must account for any dependency between the random elements of $\bm{\xi}_{x}$. Section 4 details different methods to define appropriate transformations respecting Assumption 3.
In the field of stochastic optimization, Varaiya and Wets (1989) and Dupacová (2006) introduced the idea of transforming a decision-dependent probability distribution into an independent distribution in what is called “plug-in” or “change of measure” (see Pflug 2012 for a detailed explanation) and might employ random variable transformation. Despite these initial works, no study has directly modeled endogenous uncertainty as a function $t(.)$ nor investigated appropriate transformations for random variables following classical endogenous probability distributions. Specifically, the definition of $t(.)$ remains unexplored. Moreover, no study has focused on the nature and solution of problems resulting from this transformation nor, notably, addressed their potential nonconvexity. Finally, Holzmann and Smith (2021) used random variable transformation for a specific endogenous discrete random variable, for which Section 4.1.1 highlights the similarities and differences with the approach proposed herein. Therefore, the present work proposes a general framework using functions of random variables for addressing 2-SMIPD given Assumption 3. We also investigate practical transformations for endogenous random variables following classical probability distributions, as well as the convexity of the resulting transformed program.
Based on Assumption 3, we can formulate 2-SMIPD as the transformed program
[…]
where the transformed second stage program $Q_{T}\bm{(}x,\vartheta(\omega)\bm{)}$ is defined as
$\min\{t_{1}\bm{(}x,\widetilde{q}\left(\omega\right)^{\top}\bm{)}y(\omega):t_{3}\bm{(}x,\widetilde{T}\left(\omega\right)\bm{)}x+t_{4}\bm{(}x,\widetilde{W}\left(\omega\right)\bm{)}y(\omega)\geq t_{2}\bm{(}x,\widetilde{h}\left(\omega\right)\bm{)},y(\omega)\in\mathcal{Y}\},$ (5)
and we can outline the following lemma:
Lemma 3.1
Under Assumption 3, $g(x)=\mathbb{E}_{\bm{\xi}\mid x}\left[Q\bm{(}x,\xi(\omega)\bm{)}\right]=\mathbb{E}_{\bm{\vartheta}}\left[Q_{T}\bm{(}x,\vartheta(\omega)\bm{)}\right]$ and the 2-SMIPD (Program (1)) can be reformulated into the transformed Program (4). In other words, Programs (1) and (4) have the same optimal solutions.
Thus, our methodology is established on finding an appropriate transformation function to obtain Program (5). In that way, under Assumption 3, 2-SMIPD can be transformed into a two-stage stochastic program with exogenous uncertainty. Solution procedures for the transformed Program (4) based on its properties are well-studied (see, e.g., Shapiro et al. 2021) and, thus, can be applied in this case. Program (5) has the terms $t_{1}\bm{(}x,\widetilde{q}\left(\omega\right)^{\top}\bm{)}y$, $t_{2}\bm{(}x,\widetilde{h}\left(\omega\right)\bm{)}$, $t_{3}\bm{(}x,\widetilde{T}\left(\omega\right)\bm{)}x$, and $t_{4}\bm{(}x,\widetilde{W}\left(\omega\right)\bm{)}y$, which might be nonlinear nonconvex. We then introduce the following assumption: {assumption} The continuous relaxation of the second-stage Program (5) is convex. Program (5) might need to be reformulated to convexify its continuous relaxation. Section 4 presents different cases for which Assumption 3 holds. In addition, the exogenous random vector $\bm{\vartheta}$ might have a large or infinite dimension (see, e.g., the inversion in Section 4.1). Hence, next, we describe the application of the SAA (Shapiro et al. 2021, Chap. 5) to solve Program (4), which, although intuitive, shows that the transformation does not compromise desirable SAA properties. Nonetheless, other procedures can also be employed to solve Program (4) in this case, e.g., stochastic gradient (Shapiro et al. 2021).
3.1 Sample Average Approximation
We consider $N$ independent and identically distributed samples $(\bm{\vartheta^{1}},\ldots,\bm{\vartheta^{N}})$ of $\bm{\vartheta}$ with realizations $(\vartheta^{1},\ldots,\vartheta^{N})$. The transformed SAA program is
[…]
where $\hat{g}_{N}(x)$ is the SAA estimator of $\mathbb{E}_{\bm{\vartheta}}\left[Q_{T}\bm{(}x,\vartheta(\omega)\bm{)}\right]$ defined as
$\hat{g}_{N}(x)=\frac{1}{N}\sum_{i=1}^{N}Q_{T}(x,\vartheta^{i}).$ (7)
This estimator converges point-wise with probability one (w.p.1) to $\mathbb{E}_{\bm{\vartheta}}[Q_{T}\bm{(}x,\vartheta(\omega)\bm{)}]$ when ${N\rightarrow+\infty}$. From Lemma 3.1, we have $\mathbb{E}[\hat{g}_{N}(x)]=g(x)$, so $\hat{g}_{N}(x)$ is an unbiased estimator of $g(x)$. We denote $v^{*}$, $v_{t(x,\bm{\vartheta})}^{*}$ and $v_{N}$ as the optimal values of Program (1), Program (4), and Program (6), respectively. Under some mild conditions, such as when Program (5) is convex, an MILP or when the first-stage set $\mathcal{X}$ is finite, $v_{N}$ converges to $v_{t(x,\bm{\vartheta})}^{*}$ w.p.1 as ${N\rightarrow+\infty}$ (cf. Shapiro et al. 2021, Chap. 5, and Kleywegt et al. 2002, respectively). In these cases and under Assumption 3, $v_{N}$ converges to $v^{*}$ w.p.1 as $N\rightarrow+\infty$, since $v^{*}=v_{t(x,\bm{\vartheta})}^{*}$ from Lemma 3.1. This result, and its rate of convergence (cf. (Shapiro et al. 2021)), depends on the properties of $\bm{\vartheta}$ (i.e., on the convergence of $v_{N}$ to $v_{t(x,\bm{\vartheta)}}$), which may have larger dimension and support than $\bm{\xi}$ and may require a larger number of samples.
The SAA finds a feasible first-stage solution $\overline{x}\in\mathcal{X}$. For that, we solve $M$ instances of the SAA Program (6), each with $N$ independent samples. We then evaluate the optimal solutions of every instance by using a larger sample size $N^{\prime}$ and define $\overline{x}$ as the optimal solution with the best evaluation. When $Q\bm{(}x,\xi(\omega)\bm{)}$ is finite and Program (1) has relatively complete recourse (i.e., the second stage associated with $x\in\mathcal{X}$ is feasible for almost every realization of $\bm{\xi}$), then $Q_{T}\bm{(}x,\vartheta(\omega)\bm{)}$ is also finite, and Program (4) also has relatively complete recourse, given that $\bm{\xi}_{x}=t(x,\bm{\vartheta})$ from Assumption 3. In this case, in addition to finding $\overline{x}$, we can compute a lower-bound estimator $\overline{v}_{M}^{N}$ of $v^{*}$, an upper-bound estimator $\hat{v}_{N^{\prime}}(\overline{x})$ of $v^{*}$, an over-estimator $\overline{gap}(\overline{x})$ of the optimality gap $c^{\top}\overline{x}+g(\overline{x})-v^{*}$, and unbiased estimators of the variances of $\overline{v}_{N}^{M}$, $\hat{v}_{N{{}^{\prime}}}(\overline{x})$, and $\overline{gap}(\overline{x})$, i.e., $\hat{\sigma}_{N,M}^{2}$, $\hat{\sigma}_{N{{}^{\prime}}}^{2}(\overline{x})$, and $\hat{\sigma}_{gap}^{2}(\overline{x})$, respectively. The Appendix 7 details these estimators and the SAA’s procedure.
By defining the $\bm{\xi}_{x}$ with the function $t(x,\bm{\vartheta})$, the estimators $\hat{g}_{N}(x)$ for all $x\in\mathcal{X}$ are based on the same samples of $\bm{\vartheta}$. Choosing a transformation $t(x,\bm{\vartheta})$ that creates a positive correlation between these estimators $\hat{g}_{N}(x)$ is desirable because it reduces the variance of their difference. Indeed, when comparing two first-stage decisions $x_{1}$ and $x_{2}$, the variance of the difference $\hat{g}_{N}(x_{1})-\hat{g}_{N}(x_{2})$ when both estimators use the same random samples of $\bm{\vartheta}$ is $Var\bm{(}\hat{g}_{N}(x_{1})\bm{)}+Var\bm{(}\hat{g}_{N}(x_{2})\bm{)}-2Cov\bm{(}\hat{g}_{N}(x_{1}),\hat{g}_{N}(x_{2})\bm{)}$, where $Var(\cdot)$ and $Cov(\cdot,\cdot)$ denote the variance and covariance. This is the reasoning behind the well-known efficiency-improvement technique of common random numbers (Glasserman and Yao 1992). This discussion extends to the application of any sampling technique for solving the transformed program. Section 4 further discusses this question.
4 Transformation Functions: Definition and Examples
This section presents appropriate transformations $\bm{\xi}_{x}=t(x,\bm{\vartheta})$ with $x\in\mathcal{X}$ for classical distributions in which both Assumptions 3 and 3 hold. We classify the transformation approaches according to four main categories to organize our discussion: inversion (Section 4.1), enumeration (Section 4.2), standardization (Section 4.3), and convolution (Section 4.4). Then, Section 4.5 discusses the proposed methodology’s advantages and disadvantages. We use the notation introduced in Section 2.1, and we denote $\bm{\xi}_{x}^{u}$ as the $u$-th element of the random vector $\bm{\xi}_{x}$, and $\bm{\vartheta}_{u}$ as the random variable or vector used to obtain $\bm{\xi}_{x}^{u}$. Similarly, $\xi_{x}^{u}$ and $\vartheta_{u}$ denote their realizations. Unless stated otherwise, we consider the random elements of $\bm{\xi}_{x}$ to be independent. If the random elements of $\bm{\xi}_{x}$ are independent, then the $\bm{\vartheta}_{u}$ for $u\in\mathcal{U}$ must also be independent. Thus, we can model each random element $\bm{\xi}_{x}^{u}$ separately by using its marginal probability distribution conditioned on $x$ and the $\bm{\vartheta}_{u}$. Nonetheless, in certain cases, when the elements of $\bm{\xi}_{x}$ are correlated, we can have correlated $\bm{\vartheta}_{u}$ and also model each $\bm{\xi}_{x}^{u}$ separately in the same way as in the independent case, as explained further in this section. As a result, we mainly focus on the individual transformation of an element $u$ in this section, i.e., $\bm{\xi}_{x}^{u}=t(x,\bm{\vartheta}_{u})$. We remark that most studies have focused on the independent case (see Section (2.2)), while our approach is suitable for both independent and dependent cases.
4.1 Inversion
The inversion method is commonly used to sample a random variable when the probability mass or density function is known (e.g., the inverse sampling method, cf. Robert and Casella 1999, Chap. 2). This technique transforms a random variable following the continuous uniform distribution $U(0,1)$ into the desired random variable. We define $\mathbb{F}_{\bm{\xi}^{u}}(x,\beta)(\cdot)$ to be the conditional marginal cumulative distribution function of the random variable $\bm{\xi}^{u}$ given $x$, the first-stage variables, and $\beta$, a set of remaining parameters that are not decisions of the optimization model. To apply the inverse transformation, we compute the generalized inverse of the conditional cumulative distribution:
$\mathbb{F}^{-1}_{\bm{\xi}^{u}}(x,\beta)(j)=\min\left\{\xi\in\Xi_{x}\mid\mathbb{F}_{\bm{\xi}^{u}}(x,\beta)(\xi)\geq j\right\}.$ (8)
The transformation is then defined as $\bm{\xi}_{x}^{u}=t(x,\bm{\vartheta}_{u})=\mathbb{F}^{-1}_{\bm{\xi}^{u}}(x,\beta)(\bm{\vartheta}_{u})$ (Property 2 of Assumption 3), such that $\bm{\vartheta}_{u}$ follows the $U(0,1)$ and is independent of $x$ (Property 1 of Assumption 3). The inversion method, in many cases, has the desirable property of inducing a positive correlation (see Section 3.1) between the estimators of different systems (i.e., the $\hat{g}_{N}(x)$ for $x\in\mathcal{X}$) when their samples are generated from the same sequence of uniform random numbers (Whitt 1976).
If the elements of $\bm{\xi}_{x}$ are dependent and their correlation matrix is known, then we can define correlated random elements $\bm{\vartheta}_{u}$ for $u\in\mathcal{U}$ using an appropriate copula. A copula is a multivariate cumulative distribution function that ensures the marginal distribution of each random variable $\bm{\vartheta}_{u}$ is the $U(0,1)$. Copulas are well-known for modeling dependencies (see, e.g., Nelsen 2006). Thus, as in the independent case, each $\bm{\vartheta}_{u}$ can be transformed into $\bm{\xi}_{x}^{u}$ separately using the marginal distribution of this endogenous random element (i.e., the transformation is equal for both independent and dependent cases; the difference lies in the definition of $\bm{\vartheta}_{u}$). To keep the paper self-contained, we illustrate the application of the inverse in Appendix 9 (Example 9.1) for both cases. We note that this method is limited to cases where the function $\mathbb{F}^{-1}_{\bm{\xi}^{u}}(x,\beta)(j)$ has a well-defined or approximated form. Next, we investigate the application of this technique for discrete random variables and continuous random variables according to the properties of their cumulative probability distribution.
4.1.1 Discrete Random Variables.
In this case, the inverse transformation $\bm{\xi}_{x}^{u}=\mathbb{F}^{-1}_{\xi^{u}}(x,\beta)(\bm{\vartheta}_{u})$ is simplified and consists of defining the realization $r\in\mathcal{R}_{u}$ such that
$\sum_{j=1}^{r-1}\hat{p}_{j}^{u}(x)\leq\vartheta_{u}<\sum_{j=1}^{r}\hat{p}_{j}^{u}(x),$ (9)
where the sum on the left-hand-side equals zero when $r=1$, and $\hat{p}_{j}^{u}(x)$ and $\sum_{j=1}^{r}\hat{p}_{j}^{u}(x)$ are the marginal and cumulative probability functions (see Section 2.1) of the realization $r$ of random element $u\in\mathcal{U}$. Thus, we define the variables $\theta_{u}^{r}\in\{0,1\}$, which take the value one if the realization $r\in\mathcal{R}_{u}$ respects Inequalities (9). We also have the variables $\pi_{u}^{r}\in\{0,1\}$, which equal the sum $\sum_{j=1}^{r}\theta_{u}^{j}$ for every realization $r\in\mathcal{R}_{u}$. The following constraints model Inequalities (9):
$\pi_{u}^{r}\leq\frac{1}{\vartheta_{u}}\sum_{j=1}^{r}\hat{p}_{j}^{u}(x)-\epsilon$ $\forall\ r\in\{1,\ldots,R_{u}-1\}$ (10)
$(1-\pi_{u}^{r-1})\leq\frac{1}{(1-\vartheta_{u})}\sum_{j=r}^{R_{u}}\hat{p}_{j}^{u}(x)$ $\forall\ r\in\{2,\ldots,R_{u}\}$ (11)
$\pi_{u}^{r}=\pi_{u}^{r+1}-\theta_{u}^{r+1}$ $\forall\ r\in\{1,\ldots,R_{u}-1\}$ (12)
$\sum_{r\in\mathcal{R}_{u}}\theta_{u}^{r}=1,$ (13)
where $\epsilon>0$ is a sufficiently small value used to define the strict inequality in (9). Constraints (10) and (11) ensure that the right- and left-hand side in Inequalities (9) are respected, respectively. Constraints (12) establish the connection between $\pi^{r}_{u}$ and $\theta^{r}_{u}$ so that $\theta^{r}_{u}=1$ for the right realization, where $\pi_{u}^{R_{u}}$ is set to one. Although we could model Constraints (10) and (11) with variables $\theta_{u}^{r}$, we use the variables $\pi_{u}^{r}$ to reduce the number of nonzero elements in these constraints. Constraint (13) restricts the choice of a unique realization for element $u$. Finally, $\nu_{r}^{u}$ is the value of realization $r$ of the random element $u$, so $t(x,\vartheta_{u})=\sum_{r=1}^{R_{u}}\nu_{r}^{u}\theta_{u}^{r}$. To ensure Assumption 3, the continuous relaxation of Constraints (10) and (11) must be convex. This leads to the following theorem:
Theorem 4.1
If all the functions $\hat{p}^{u}_{r}(x)$ for $u\in\mathcal{U}$ and $r\in\mathcal{R}_{u}$ are either convex or concave over $\left[0,1\right]^{n_{1}}$, then Constraints (10) and (11) have a convex continuous relaxation.
The Appendix 8 presents the proof of Theorem 4.1. Different studies investigate problems with convex or concave $p_{r}^{u}(x)$ (see Section 2.2). Despite this desirable property, the original program is usually nonconvex. In contrast, using the inverse transformation function results in only linear or convex constraints. In addition, even if the continuous relaxation of the $\hat{p}_{r}^{u}(x)$ is neither concave nor convex, we can convexify Constraints (10) and (11) in certain cases (this analysis is problem-specific). Nonetheless, in addition to relying on the probability function $\hat{p}^{u}_{r}(x)$ to ensure convexity, the number of integer variables and constraints increases polynomially with the number of random parameters $U$ and their realizations $R_{u}$, which impacts the problem’s complexity and highlights the computational challenges of handling endogenous uncertainty.
To the best of our knowledge, only Holzmann and Smith (2021) used inversion to model a specific discrete endogenous random variable in 2-SMIPD. In contrast, we describe a general framework adapted to various functions $\hat{p}_{r}^{u}(x)$. The set of Constraints (10)–(13) also has a tighter continuous relaxation feasibility set than that employed by Holzmann and Smith (2021) (see Appendix 10).
In certain cases, such as when $\bm{\xi}_{x}^{u}$ has only two possible realizations and $p_{r}^{u}(x)$ is a linear function, simpler models can be designed. This is the case of the Bernoulli distribution, described next.
Bernoulli Distribution.
We consider that $\bm{\xi}_{x}^{u}\sim Bern(\phi)$, with the probability of success $\phi\in[0,1]$ being a subvector of $x$. The inverse is such that $\xi_{x}^{u}=1$ if $\vartheta_{u}<\phi$, and $\xi_{x}^{u}=0$ otherwise, with $\bm{\vartheta}_{u}\sim U(0,1)$. We model this with the variables $\upsilon_{u}\in\{0,1\}$, i.e., $t(x,\vartheta_{u})=\upsilon_{u}$, and the constraints
$\{\vartheta_{u}\upsilon_{u}\leq\phi-\epsilon\;;\;\upsilon_{u}\geq\phi-\vartheta_{u}\},$ (14)
where $\epsilon>0$ is again a sufficiently small value. Constraints (14) guarantee that realization 00 and $1$ are chosen if $\vartheta_{u}$ is superior to $\phi$ and if $\vartheta_{u}$ is inferior or equal to $\phi$, respectively.
For both transformations (general and specific to Bernoulli), if only $h(\omega)$ is uncertain, then Assumption 3 is respected. Otherwise, linearization techniques, such as McCormick, can be applied to linearize the bilinear terms in $t_{1}\bm{(}x,\widetilde{q}(\omega)^{\top}\bm{)}y$, $t_{3}\bm{(}x,\widetilde{T}(\omega)\bm{)}x$, and $t_{4}\bm{(}x,\widetilde{W}(\omega)\bm{)}y$, so Assumption 3 holds. In this case, the problem’s complexity increases further, and the transformation is restricted by the dimensions of variables $x$ and $y$, due to the inclusion of extra variables and constraints when applying linearization techniques.
4.1.2 Continuous Random Variables.
In general, the transformation $\xi_{x}^{u}=\mathbb{F}^{-1}_{\xi^{u}}(x,\beta)(\vartheta_{u})$ cannot be simplified for continuous random variables. Nevertheless, when, for fixed $x$ and $\beta$, the conditional cumulative distribution is strictly increasing and continuous, the generalized inverse is redundant because $\mathbb{F}^{-1}_{\bm{\xi}^{u}}(x,\beta)(\cdot)$ is the unique inverse function. This property is commonly respected for well-behaved distributions and leads to a simpler transformation. Two applications of the inverse for some well-known continuous distributions respecting this property are listed below.
Uniform distribution.
If $\bm{\xi}^{u}\sim U(a,b)$ and both parameters $a\in\mathbb{R}$ and $b\in\mathbb{R}$ are first-stage variables, then we can use the transformation function $t(a,b,\bm{\vartheta}_{u})=a+(b-a)\bm{\vartheta}_{u}$.
Exponential distribution.
If $\bm{\xi}^{u}\sim\exp(\lambda)$ with $\lambda>0$ a first-stage decision, then the transformation is defined as $t(\lambda,\bm{\vartheta}_{u})=-{\ln(1-\bm{\vartheta}_{u})}/{\lambda}$. This function is concave for $\lambda>0$.
Both distributions are part of the special case where the unique inverse $\mathbb{F}^{-1}_{\bm{\xi}^{u}}(x,\beta)(\cdot)$ is convex or concave on $x\in\mathcal{X}$, i.e., the inverse function is convex or concave on the parameters of the distribution. Other distributions that fall into this special case are the Cauchy and logistic distributions. Additionally, there are distributions for which $\mathbb{F}^{-1}_{\bm{\xi}^{u}}(x,\beta)(\cdot)$ has no closed-form expression, but good numerical approximations are available, e.g. student, chi-square, and normal distributions. Nevertheless, transformations based on other techniques, which lead to simpler formulations, can be used for certain distributions (see Section 4.3 for the normal distribution). The analysis of the special class of distributions introduced before leads to the following theorem:
Theorem 4.2
Assumption 3 holds if the following criteria are met: (1) the cost vector $q$, technology matrix $T$ and recourse matrix $W$ are fixed; (2) $\mathbb{F}_{\bm{\xi}^{u}}(x,\beta)(\cdot)$for fixed $x$ and $\beta$ is continuous and strictly increasing; and (3) the transformation $t(x,\bm{\vartheta}_{u})=\mathbb{F}^{-1}_{\bm{\xi}^{u}}(x,\beta)(\cdot)$ is a convex function on $x\in\mathcal{X}$, or concave if the recourse constraint in Program (5) is an inferior or equal to inequality.
The Appendix 8 presents the proof of Theorem 4.2. If Criterion (1) does not hold, the products $t_{1}\bm{(}x,\widetilde{q}(\omega)^{\top}\bm{)}y$, $t_{3}\bm{(}x,\widetilde{T}(\omega)\bm{)}x$ and $t_{4}\bm{(}x,\widetilde{W}(\omega)\bm{)}y$ may be nonconvex even if Criterion (3) holds. Thus, a problem-specific convexity analysis is required to confirm if Assumption 3 is satisfied. This evidences the complexity added by endogenous uncertainty, as the transformation can, in some cases, produce a nonconvex nonlinear transformed program with exogenous uncertainty, which is theoretically intractable.
4.2 Enumeration
For a random element $\bm{\xi}_{x}^{u}$ with $u\in\mathcal{U}$, a simple transformation can be employed if the set of alternative marginal distributions $\mathcal{D}_{u}$ of element $u$ is sufficiently small to allow enumeration. This corresponds to the “discrete selection of distributions” (see Section 2.1). We have $\bm{\vartheta}_{u}=\{\bm{\vartheta}_{ud}\}_{d\in\mathcal{D}_{u}}$, where $\bm{\vartheta}_{ud}$ is a random variable following distribution $d\in\mathcal{D}_{u}$. The $\bm{\vartheta}_{u}$ is independent of $x$ (Property 1 of Assumption 3). We thus have ${t(x,\bm{\vartheta}_{u})=\sum_{d=1}^{D_{u}}\bm{\vartheta}_{ud}x_{u}^{d}}$ (Property 2 of Assumption 3), with $x_{u}^{d}$ as defined in Section 2.1. This transformation can be applied for $\bm{\xi}_{x}$ with independent or correlated random elements by also using copulas. To keep the paper self-contained, we illustrate this in Appendix 9, Example 9.2, and also present the case where the distribution is selected for the entire random vector $\bm{\xi}_{x}$ in Example 9.3. If only $h(\omega)$ is uncertain, then Assumption 3 is directly satisfied. Otherwise, bilinear terms may appear due to $t_{1}\bm{(}x,\widetilde{q}(\omega)^{\top}\bm{)}y$, $t_{3}\bm{(}x,\widetilde{T}(\omega)\bm{)}x$ and $t_{4}\bm{(}x,\widetilde{W}(\omega)\bm{)}y$. In such cases, auxiliary linearizations such as the McCormick envelope can be applied so Assumption 3 is respected. However, this approach is limited not only by the size of $\mathcal{D}_{u}$ but also by the dimensions of $y$ or $x$, as it increases the number of variables, leading to computational challenges.
4.3 Standardization
A location-scale family of distributions has distributions with location $\mu$ and scale $\sigma$ parameters (Casella and Berger 2024, Chap. 3.5). The distributions of this family can be derived from a standard family distribution member with location and scale parameters equal to one and zero, respectively. If $\bm{\xi}^{u}$ follows a location-scale family distribution with parameters $\mu$ and $\sigma$ as first-stage decisions, then $t(\mu,\sigma,\bm{\vartheta}_{u})=\mu+\sigma\bm{\vartheta}_{u}$ is valid (Property 2 of Assumption 3). The $\bm{\vartheta}_{u}$ follows the standard family distribution member, which is independent of the values of $\mu$ and $\sigma$ (Property 1 of Assumption 3). If only the vector $h(\omega)$ has random elements, then Assumption 3 holds. Otherwise, a specific analysis of the terms $t_{1}(x,\widetilde{q}(\omega)^{\top})y$, $t_{3}(x,\widetilde{T}(\omega))x$ and $t_{4}(x,\widetilde{W}(\omega))y$ should be conducted. Thus, similar to the inversion, the transformed program may also be nonlinear and nonconvex, resulting in computational challenges. Many distributions belong to the location-scale family, including normal, generalized extreme value, exponential, and logistic distributions. Next, we exemplify this method for two of these distributions.
Normal distribution.
If $\bm{\xi}^{u}\sim N(\mu,\sigma^{2})$ and $\mu$ and $\sigma$ are first-stage decisions, then we can apply the proposed transformation with $\bm{\vartheta}_{u}$ following the standard normal $N(0,1)$.
Generalized extreme value distribution.
If $\bm{\xi}^{u}\sim GEV(\mu,\sigma,\beta)$ with $\mu$ and $\sigma$ first-stage decisions and $\beta$ a fixed parameter, then we can apply the proposed transformation with $\bm{\vartheta}_{u}\sim GEV(0,1,\beta)$.
4.4 Convolution
Various random variables that follow well-known distributions can be expressed as the sum of other independent random variables. Notable examples include the binomial distribution, which is the sum of Bernoulli random variables, and the hypoexponential distribution, which is the sum of exponential random variables. Suppose that $\bm{\xi}_{x}^{u}$ is equal to the sum $\bm{\zeta}_{x}^{u1}+\bm{\zeta}_{x}^{u2}+\ldots+\bm{\zeta}_{x}^{un}$ of independent random variables $\bm{\zeta}_{x}^{uj}$’s following a specified endogenous distribution. In this case, the probability distribution of $\bm{\xi}_{x}^{u}$ is the convolution of the marginal distributions of the $\bm{\zeta}_{x}^{uj}$’s (Hogg et al. 2013, Chap. 2). We can define a transformation $\bm{\zeta}_{x}^{uj}=\hat{t}_{j}(x,\bm{\vartheta}_{uj})$ for every endogenous random variable $\bm{\zeta}_{x}^{uj}$, $j\in\{1,\ldots,n\}$, independently. Thus, $t(x,\bm{\vartheta}_{u})=\sum_{j=1}^{n}\hat{t}_{j}(x,\bm{\vartheta}_{uj})$, with $\bm{\vartheta}_{u}=\{\bm{\vartheta}_{uj}\}_{j=1}^{n}$. The complexity of the problem depends on the definition of each transformation $\hat{t}_{j}(x,\bm{\vartheta}_{uj})$ and the number of components $n$, thus limiting the application of this method by the size of $n$. Next, we demonstrate this transformation for the binomial distribution. We also provide an example of dependent random variables for which this methodology can be applied in Appendix 9 (Example 9.4).
Binomial distribution.
We consider that $\bm{\xi}_{x}^{u}\sim Bin(n,\phi)$, where the probability of success of one trial $\phi\in[0,1]$ is a subvector of $x$, and the number of trials $n$ is a parameter. We have $\bm{\xi}_{x}^{u}=\bm{\zeta}_{x}^{u1}+\bm{\zeta}_{x}^{u2}+\ldots+\bm{\zeta}_{x}^{un}$, where the $\bm{\zeta}_{x}^{uj}$’s are endogenous and follow the $Bern(\phi)$ independently. We define the $\bm{\zeta}_{x}^{uj}$’s with the inversion for the Bernoulli distribution of Section 4.1.1. We have ${\bm{\vartheta}_{uj}\sim U(0,1)}$ independently for $j\in\{1,\ldots,n\}$. We set the variables $\upsilon_{ur}\in\{0,1\}$ equal to one if realization $\vartheta_{ur}<\phi$ and zero otherwise, and include Constraints (14) for all $\bm{\zeta}_{x}^{uj}$’s. Thus, $t(x,\vartheta_{u})=\sum_{j=1}^{n}\upsilon_{uj}$. If only $h(\omega)$ is uncertain, then Assumption 3 is respected. Otherwise, the bilinear terms in $t_{1}(x,\widetilde{q}(\omega)^{\top})y$, $t_{3}(x,\widetilde{T}(\omega))x$, and $t_{4}(x,\widetilde{W}(\omega))y$ can be linearized by using the McCormick envelope (Assumption 3 holds). In this case, the application of this technique is limited to the dimensions of $y$ and $x$ due to the increase in the number of variables. Furthermore, inversion (see Section 4.1.1) may be used in this case because the binomial probability of every realization is a convex function on the probability of success $\phi$ (i.e., we can apply Theorem 4.1). Both approaches have a similar number of variables and constraints and include big-M constraints. However, inversion leads to convex (potentially nonlinear) constraints, whereas convolution leads to linear constraints.
4.5 Advantages and Disadvantages
The transformed Program (4), a two-stage program with exogenous uncertainty, offers several advantages: (i) Solution procedures for such programs are extensively researched and can be applied based on their specific properties. (ii) Large-scale decomposition techniques, such as the L-shaped algorithm, are applicable depending on the characteristics of the transformed second-stage Program (5). (iii) Sampling techniques become straightforward, enabling the treatment of problems with continuous endogenous random variables or discrete ones with many realizations. In contrast, using sampling to solve the original program with endogenous uncertainty is challenging, as the distribution depends on first-stage decisions. Besides, our method is versatile as various techniques can be employed to define the transformation function $t(x,\bm{\vartheta})$. As a result, multiple transformations may exist to model the same endogenous random vector. For instance, both the inverse and convolution methods can be applied to model the binomial random variable. The choice of a transformation should aim to achieve a transformed program with desirable properties. Furthermore, as demonstrated earlier in this section, in many cases, a transformation function leading to a mixed-integer linear or convex transformed Program (4) exists, even if the original Program (1) is nonlinear nonconvex. This is particularly the case for problems where only the vector $h(\omega)$ is uncertain, as the transformation $t_{2}\bm{(}x,\widetilde{h}\left(\omega\right)\bm{)}$ does not multiply the variables $x$ or $y$ in Program (5). For the discrete selection of distributions where only $h(\omega)$ is uncertain, the original program is an MILP obtained by linearizing the objective function with many variables and big-M constraints (see (Bhuiyan et al. 2020)). However, we propose a transformation based on the enumeration of distributions that reduces the need for additional variables or constraints in this case.
Nevertheless, certain transformations, such as enumeration and the inverse transformation for discrete random variables, may introduce in the formulation many constraints and variables, which are often integer when dealing with discrete distributions. Hence, we explore the impact of such transformations in our experimental results. For problems where the cost vector $q(\omega)$, technology matrix $T(\omega)$, and/or recourse matrix $W(\omega)$ are uncertain, obtaining a transformed program with a convex relaxation is more challenging due to the terms $t_{1}(x,\widetilde{q}(\omega)^{\top})y$, $t_{3}(x,\widetilde{T}(\omega))x$, and $t_{4}(x,\widetilde{W}(\omega))y$. For discrete distributions, introducing integer variables to define the transformations enables the use of McCormick linearization as a last resort. However, this approach is constrained by the dimensions of $x$, $y$, and the number of uncertain parameters. Additionally, the dimension and support of the exogenous random vector $\bm{\vartheta}$ are typically larger than those of the endogenous random vector $\bm{\xi}_{x}$. The convergence properties of sampling techniques depend on the support of $\bm{\vartheta}$, potentially requiring more samples. However, leveraging the same random samples of $\bm{\vartheta}$ for the SAA estimators $\hat{g}_{N}(x)$ for all $x\in\mathcal{X}$ to induce correlation between them is consistent with the efficiency-improvement technique of common random numbers (see discussion in Section 3.1).
5 Case Study
Pre-disaster or pre-interdiction problems with endogenous uncertainty have been investigated in the context of facility protection (Medal et al. 2016, Bhuiyan et al. 2020, Li et al. 2021, 2022) and network design protection (Viswanath et al. 2004, Flach 2010, Peeta et al. 2010, Du and Peeta 2014, Prestwich et al. 2014). To the best of our knowledge, these applications are the most recurrent in previous research. Thus, this section showcases our methodology by applying it to the NDFPP, which was introduced by Bhuiyan et al. (2020)Bhuiyan et al. (2020) defined the endogenous random variables through a discrete selection of distributions. In contrast, we explore this and three other decision-dependent probability mass or density functions for the random variables. This section introduces our transformed program for the NDFPP, which has different versions according to the definition of the endogenous distribution.
5.1 Network Design and Facility Protection Problem
The NDFPP is a 2-SMIPD with complete recourse. First-stage decisions determine facility protection allocations and the network structure. Next, the scenario (natural disruptions) is realized, affecting facility capacity levels, and second-stage decisions establish an optimal flow to satisfy client demands. Facility capacity levels are mutually independent endogenous random variables.
We consider a set $\mathcal{D}_{0}$ of natural disruption events, representing different types of natural disasters centered at specific locations. Facilities experience varying disruption intensities based on their distance from the disruption center $d\in\mathcal{D}_{0}$, with intensity levels in the set =$\mathcal{L}=\{1,\ldots,L\}$. We define the set of possible events $\mathcal{D}=\mathcal{D}_{0}\cup\{nd\}$, which includes the no-natural-disruption event $nd$. The network is represented by the undirected graph $G=(\mathcal{N}_{\mathcal{T}},\mathcal{E})$, where $\mathcal{N}_{\mathcal{T}}=\mathcal{N_{F}}\cup\mathcal{N_{C}}\cup\{m\}$ consists of facilities ($\mathcal{N_{F}}$), clients ($\mathcal{N_{C}}$), and a dummy facility ($m$). The set $\mathcal{N}_{\mathcal{T}}^{0}=\mathcal{N}_{\mathcal{T}}\backslash\{m\}$ excludes the dummy facility. We also specify a directed graph $G_{o}=(\mathcal{N}_{\mathcal{T}},\mathcal{A})$, including arcs $(i,j)$ and $(j,i)$ for each edge $[i,j]\in\mathcal{E}$. The edge corresponding to arc $(i,j)\in\mathcal{A}$ is denoted as $e(i,j)$. To represent unsatisfied demand, we include arcs $(m,n)$ for every client node $n\in\mathcal{N_{C}}$ in $\mathcal{A}$, with the auxiliary set of arcs arriving at the dummy facility defined as $\mathcal{A}_{m}=\{(m,n)\ \forall\ n\in\mathcal{N_{C}}\}$. The set $\mathcal{P}=\{1,\ldots,P\}$ contains available protection levels for each facility, where larger values indicate better protection. Finally, $\mathcal{S}$ is the set of scenarios, each composed of an event and a realization of the exogenous random vector. The function $\hat{d}(s)$ returns the event $d\in\mathcal{D}$ for scenario $s\in\mathcal{S}$. The realization of the exogenous random vector in scenario $s$ is denoted $\vartheta^{s}$. The definition of this random vector depends on the endogenous distribution and transformation function, which we define later. Moreover, protecting a facility $f\in\mathcal{N_{F}}$ with protection level $p\in\mathcal{P}$ costs $c_{f}^{p}$, and opening an edge $e\in\mathcal{E}$ costs $c_{e}$. Each client node $n\in\mathcal{N_{C}}$ has a demand $b_{n}$. The budget for edge opening and facility protection is denoted by $C$. The per-unit-flow transportation cost for arc $(i,j)\in\mathcal{A}$ is $q_{ij}$. Moreover, the cost for arcs $\mathcal{A}_{m}$ leaving the dummy facility is $a\cdot\max_{(i,j)\in\mathcal{A}\backslash\mathcal{A}_{m}}q_{ij}$, where $a$ is the penalty per unit of unmet demand. The big-M value for edge capacity is $B=\sum_{n\in\mathcal{N_{C}}}b_{n}$. For the first stage, we define the binary variables $x_{f}^{p}$, which are equal to one if protection level $p\in\mathcal{P}$ is allocated to facility $f\in\mathcal{N_{F}}$ and zero otherwise. We also include the binary variables $z_{e}$, which are equal to one if edge $e\in\mathcal{E}$ is installed and zero otherwise. For the second stage, we specify the variables $y_{ij}^{s}$, which represent the flow through arc $(i,j)\in\mathcal{A}$ in scenario $s\in\mathcal{S}$. The endogenous capacity of facilities, $\xi_{x}^{fs}$ for facility $f\in\mathcal{N_{F}}$ in scenario $s\in\mathcal{S}$, depends on the $x$ variables. We summarize this notation in Appendix 15. The transformed program for the NDFPP is
$\min\mbox{: }$ $\mathbb{E}_{\bm{\vartheta}}[Q_{T}(x,\bm{\vartheta})]$ (15a)
s.t. $\sum_{p\in\mathcal{P}}x_{f}^{p}=1$ $\quad\forall\ f\in\mathcal{N_{F}}$ (15b)
$\sum_{f\in\mathcal{N_{F}}}\sum_{p\in\mathcal{P}}c_{f}^{p}x_{f}^{p}+\sum_{e\in\mathcal{E}}c_{e}z_{e}\leq C$ (15c)
$x_{f}^{p}\in\{0,1\}$ $\quad\forall\ f\in\mathcal{N_{F}},\quad p\in\mathcal{P}$ (15d)
$z_{e}\in\{0,1\}$ $\quad\forall\ e\in\mathcal{E},$ (15e)
where the transformed second-stage program $Q_{T}(x,\vartheta_{s})$ for realization $\vartheta_{s}$ of $\bm{\vartheta}$ is
$\min\mbox{: }$ $\sum_{(i,j)\in\mathcal{A}}q_{ij}y_{ij}^{s}$ (16a)
$\sum_{j\in\mathcal{N_{T}}}y_{jn}^{s}-\sum_{j\in\mathcal{N_{T}}}y_{nj}^{s}=b_{n}$ $\quad\forall\ n\in\mathcal{N_{C}}$ (16b)
$\sum_{j\in\mathcal{N_{T}}}y_{fj}^{s}-\sum_{j\in\mathcal{N_{T}}}y_{jf}^{s}\leq\xi_{x}^{fs}$ $\quad\forall\ f\in\mathcal{N_{F}}$ (16c)
$y_{ij}^{s}\leq Bz_{e(i,j)}$ $\quad\forall\ (i,j)\in\mathcal{A}\backslash\mathcal{A}_{m}$ (16d)
$\xi_{x}^{fs}=t(x,\vartheta^{s})$ $\quad\forall\ f\in\mathcal{N_{F}}$ (16e)
$y_{ij}^{s}\geq 0$ $\quad\forall\ (i,j)\in\mathcal{A}.$ (16f)
In the first stage, the objective function (15a) minimizes the expected value of the transformed second-stage program. We also have Constraints (15b) and (15c), which specify the selection of one level of protection for each facility and ensure that the cost budget for investing in facility protection and edge installation, respectively, is respected. The objective function (16a) of the second stage minimizes the average cost of transporting goods from facilities to clients. In addition, the second stage contains Constraints (16b)–(16e). These ensure flow conservation at client nodes (Constraints (16b)), the use only of available capacity (Constraints (16c)), the existence of flow only in installed arcs (Constraints (16d)), and a suitable definition of the endogenous realizations $\xi_{x}^{fs}$ (Constraints (16e)). We note that, following the notation in Program (1), only $h(\omega)$ is uncertain in this problem. We show the transformed SAA Program (6) of the NDFPP in Appendix 11.
In the following sections, we describe the various problem variants based on the decision-dependent distribution of $\xi_{x}^{fs}$. Specifically, we detail Constraints (16e) and define $\vartheta^{s}$. For discrete endogenous distributions, we assume that the post-disruption capacity of facilities can take on $|\mathcal{W}|$ possible values, where $\mathcal{W}=\{0,\ldots,W\}$ is an ordered index set. Here, level 0 represents no capacity, and level $W$ represents maximum capacity. The parameter $\nu_{w}$ for $w\in\mathcal{W}$ denotes the capacity at level $w$, with $\nu_{w+1}=\nu_{w}+\overline{\nu}$, where $\overline{\nu}$ is the increase in capacity per unit increase in capacity level.
5.1.1 Discrete Selection of Distributions.
First, we consider the discrete selection of distributions introduced by Bhuiyan et al. (2020). If a protection level $p\in\mathcal{P}$ is installed at facility $f\in\mathcal{N_{F}}$, then its capacity level is considered to follow $Bin(W,\overline{\phi}_{f}^{dp})$ when event $d\in\mathcal{D}$ is realized. The parameter $\overline{\phi}_{f}^{dp}$ is the probability of successfully increasing the capacity level by one. We follow the enumration in Section 4.2. For facility $f\in\mathcal{N_{F}}$, protection $p\in\mathcal{P}$, and scenario $s\in\mathcal{S}$, we define $\vartheta^{s}_{fp}$ as the value of the realization of the capacity level following the $Bin(W,\overline{\phi}_{f}^{dp})$. In this case, no extra variables are needed. We have $\xi_{x}^{fs}=\sum_{p\in\mathcal{P}}\vartheta^{s}_{fp}x_{f}^{p}$ for every facility $f\in\mathcal{N_{F}}$, which model Constraints (16e). Finally, this variant of the considered problem is denoted NDFPP-Selection.
5.1.2 Binomial Distribution.
We assume that the network-wide protection allocation affects the capacity level distribution of all facilities. Thus, we aim to represent situations in which facilities collaborate by sharing resources, situations in which facilities intercommunicate with transmission lines (e.g., power systems), or situations where failures propagate in the network (e.g., the spread of diseases). We consider that the capacity level of a facility $f\in\mathcal{N_{F}}$ follows $Bin(W,{\phi}_{f}^{d})$ when event $d\in\mathcal{D}$ occurs, where ${\phi}_{f}^{d}\in[0,1]$ is a variable computed as:
${\phi}_{f}^{d}=\frac{1}{\rho}\sum_{i\in\mathcal{N_{F}}}\sum_{p\in\mathcal{P}}\widetilde{\phi}_{fi}^{dp}x_{i}^{p}\ \quad\forall\ d\in\mathcal{D},\quad f\in\mathcal{N_{F}},$ (17)
where the parameter $\widetilde{\phi}_{fi}^{dp}$ is the impact on the capacity level at facility $f\in\mathcal{N_{F}}$ of installing protection level $p\in\mathcal{P}$ at facility $i\in\mathcal{N_{F}}$ when event $d$ occurs, and $\rho$ is a normalization parameter to guarantee that ${\phi}_{f}^{d}\in[0,1]$. Hence, we employ the convolution technique for the Binomial distribution in Section 4.4; that is, $\vartheta_{fw}^{s}$ is the realization following the $U(0,1)$ for scenario $s\in\mathcal{S}$, facility $f\in\mathcal{N_{F}}$, and capacity level $w\in\{1,\ldots,W\}$. The variables $\upsilon^{s}_{fw}\in\{0,1\}$ are equal to one if $\vartheta_{fw}^{s}<{\phi}^{\hat{d}(s)}_{f}$ and zero otherwise for $f\in\mathcal{N_{F}}$ and $w\in\{1,\ldots,W\}$. Finally, we adapt Constraints (14) to guarantee the adequate definition of variables $\upsilon^{s}_{fw}$ (see Appendix 12). Thus, we determine the capacity value in every scenario as $\xi_{x}^{fs}=\overline{\nu}\left(\sum_{w=1}^{W}\upsilon_{fw}^{s}\right)$ for every facility $f\in\mathcal{N_{F}}$, which are equivalent to Constraints (16e). This variant of the considered problem is denoted NDFPP-Binomial. Note that the DE-OP of this variant is nonlinear nonconvex (see Appendix 13).
5.1.3 Discrete Distribution.
We continue to assume that network-wide protection influences the capacity levels of all facilities. We consider that the random variables follow a discrete distribution inspired by the multinomial logit models in discrete choice theory (Bierlaire 1998). Therefore, the probability $\hat{p}_{df}^{w}(u)$ of a facility $f\in\mathcal{N_{F}}$ having capacity level $w\in\mathcal{W}$ after the occurrence of an event $d\in\mathcal{D}$ is defined as
$\hat{p}_{df}^{w}(u)=\frac{u_{df}^{w}}{\sum_{j\in\mathcal{W}}u_{df}^{j}}\ \quad\forall\ d\in\mathcal{D},\quad f\in\mathcal{N_{F}},\quad w\in\mathcal{W},$ (18)
where $u_{df}^{w}\geq 0$ are auxiliary variables expressing the likelihood of facility $f\in\mathcal{N_{F}}$ having capacity level $w\in W$ when event $d\in\mathcal{D}$ occurs. These variables are
$u_{df}^{w}=\sum_{i\in\mathcal{N_{F}}}\sum_{p\in\mathcal{P}}\widetilde{u}_{fi}^{dpw}x_{i}^{p}\ \quad\forall\ d\in\mathcal{D},\quad f\in\mathcal{N_{F}},\quad w\in\mathcal{W},$ (19)
where the parameter $\widetilde{u}_{fi}^{dpw}$ is the impact of installing protection $p\in\mathcal{P}$ at facility $i\in\mathcal{N}_{\mathcal{F}}$ on the likelihood of capacity level $w\in\mathcal{W}$ at facility $f\in\mathcal{N_{F}}$ when event $d\in\mathcal{D}$ occurs. Thus, we adapt the inverse transformation in Section 4.1.1. For each facility $f\in\mathcal{N_{F}}$ and scenario $s\in\mathcal{S}$, the realization $\vartheta_{f}^{s}$ follows the $U(0,1)$. The variables $\theta_{fw}^{s}\in\{0,1\}$ equal one if the capacity level $w\in\mathcal{W}$ is selected for facility $f\in\mathcal{N_{F}}$ in scenario $s\in\mathcal{S}$ and zero otherwise. We also define the variables $\pi_{fw}^{s}\in\{0,1\}$, which are equal to $\sum_{j=0}^{w}\theta_{fj}^{s}$, with $\pi_{fW}^{s}$ set to one. We adapt the Constraints (10)–(13) for the NDFPP to ensure the correct inverse transformation of $\vartheta_{f}^{s}$ according to the probability distribution defined by Equation (18). The obtained constraints initially contain bilinear terms involving binary variables, and thus, to satisfy Assumption 3, we apply the exact McCormick envelope. This adaptation and linearization procedures are detailed in Appendix 12. As a result, we have $\xi_{x}^{fs}=\sum_{w\in\mathcal{W}}\nu_{w}\theta_{fw}^{s}$ for every facility $f\in\mathcal{N_{F}}$, which model Constraints (16e). This variant is denoted NDFPP-Discrete. Finally, the DE-OP in this case is also nonlinear nonconvex (see Appendix 13).
5.1.4 Normal Distribution.
We also assume that network-wide protection affects the facilities’ capacity level. We consider that the capacity level of facility $f\in\mathcal{N_{F}}$ follows $N(\mu_{f}^{d},{(\widetilde{\sigma}_{f}^{d})}^{2})$ when event $d\in\mathcal{D}$ occurs. The standard deviation $\widetilde{\sigma}_{f}^{d}\geq 0$ is a parameter and the mean $\mu_{f}^{d}\geq 0$ is a decision variable such that
$\mu_{f}^{d}=\frac{1}{\rho}\sum_{i\in\mathcal{N_{F}}}\sum_{p\in\mathcal{P}}\widetilde{\mu}_{fi}^{dp}x_{i}^{p}\ \quad\forall\ d\in\mathcal{D},\quad f\in\mathcal{N_{F}}.$ (20)
The parameter $\widetilde{\mu}_{fi}^{dp}$ is the impact of installing protection $p\in\mathcal{P}$ at facility $i\in\mathcal{N}_{\mathcal{F}}$ on the mean of the capacity level of facility $f\in\mathcal{N}_{\mathcal{F}}$ when event $d\in\mathcal{D}$ occurs, and $\rho$ is a normalization value. We ensure that the capacity is always positive, i.e., $\mu_{f}^{d}\geq 4\widetilde{\sigma}_{f}^{d}$ for all feasible decisions. We use the standardization method in Section 4.3, so the realization $\vartheta_{f}^{s}$ follows $N(0,1)$ for each facility $f\in\mathcal{N_{F}}$ and scenario $s\in\mathcal{S}$. Constraints (16e) are modeled as $\xi^{fs}_{x}=\mu_{f}^{\hat{d}(s)}+\vartheta_{f}^{s}\sigma_{f}^{\hat{d}(s)}$ for every $f\in\mathcal{N_{F}}$. This variant is denoted NDFPP-Normal, and its DE-OP is nonlinear nonconvex (see Appendix 13).
5.2 Experimental Results
This section validates the application of our methodology for addressing 2-SMIPD. First, Section 5.2.1 summarizes the instance generation. Section 5.2.2 then compares the SAA applied to the transformed program (SAA-TP) and the DE-OP for NDFPP-Selection, specifically, the MILP reformulation obtained by auxiliary linearizations in Bhuiyan et al. (2020). The DE-OP for the NDFPP-Binomial, NDFPP-Discrete, and NDFPP-Normal is nonlinear nonconvex (see Section 5). Thus, Section 5.2.3 only explores the performance of the estimators obtained by the SAA-TP. Finally, we investigate the value of the stochastic solution (VSS) of the NDFPP in Section 5.2.4. Note that different methods can be used to solve our obtained two-stage program with exogenous uncertainty. We apply the SAA, given its simplicity and desirable properties (see Section 3.1).
The instances of the transformed SAA program, the DE-OP, the EV problem, and the expected result of the EV solution (EEV) for computing the VSS were coded in C++ and solved directly with Gurobi 10.0.3. All experiments were processed on the Beluga cluster of the Compute Canada network on an Intel Gold 6148 Skylake @ 2.4 GHz with a maximum of 100 GB. We limited Gurobi to using one thread. The details on the application of the SAA are given in Appendix 7. All instances, results, and code are publicly available on GitHub (https://github.com/mariabazotte/saa_endog).
5.2.1 Instance Generation.
We follow a procedure similar to that of Bhuiyan et al. (2020) for network, transportation costs, and client demand generation, which is detailed in Appendix 16 and available on the same GitHub project. We consider five networks, with $|\mathcal{N}_{\mathcal{T}}^{0}|=15$, 25, 30, 39, and 48, and assume the values 4 and 5 for $|\mathcal{N}_{\mathcal{F}}|$. The protection costs $c_{f}^{p}$, which are randomly generated, and the edge-opening costs $c_{e}$ are defined similarly as in Bhuiyan et al. (2020) (see Appendix 16). The budget $C$ is a percentage of the total cost with maximum protection, $C=0.5(\sum_{n\in\mathcal{N_{F}}}c_{n}^{P}+\sum_{e\in\mathcal{E}}c_{e})$, making it adjustable to different instance sizes. We consider four protection levels ($P=4$), and define the maximum facility capacity as $\nu_{\text{max}}=|{\sum_{n\in\mathcal{N_{C}}}b_{n}}/{0.9|\mathcal{N_{F}}|}|$. Thus, the capacity increase $\overline{\nu}$ is ${\nu_{\text{max}}}/{W}$ with capacity values $\nu_{w}=w\overline{\nu}$ for $w\in\mathcal{W}$. We use $W=2,3,$ and $4$. As in Bhuiyan et al. (2020), we consider three disruption events with distinct centers of occurrence ($|\mathcal{D}_{0}|=3$), and we assume three disruption intensity levels ($L=3$). We use the procedure in Bhuiyan et al. (2020) to define the intensity level $\hat{l}(f,d)$ in $\mathcal{L}$ at which facility $f\in\mathcal{N_{F}}$ is impacted by disruption $d\in\mathcal{D}_{0}$ (see Appendix 16).
Analogously to Bhuiyan et al. (2020), we define the probability of successfully increasing the capacity level by one, $\overline{\phi}_{f}^{dp}$, as $({0.95\times c_{f}^{p}}/{c_{f}^{P}})^{{\hat{l}(f,d)}/{L}}$, with a 5% risk of protection malfunction for each capacity level increase. For the $nd$ event, $\overline{\phi}_{f}^{nd,p}$ is defined as $\eta({0.95\times c_{f}^{p}}/{c_{f}^{P}})+(1-\eta)$, where $\eta=0.7$ reflects the influence of protection decisions on capacity level uncertainty on this event. Unlike Bhuiyan et al. (2020), we assume that facilities are affected by the $nd$ event and low-intensity disruption events, which results in instances with a larger number of scenarios. For the NDFPP-Binomial and NDFPP-Discrete, we set $\widetilde{\phi}_{fi}^{dp}$ and $\widetilde{u}_{fi}^{dpw}$ to $\overline{\phi}_{i}^{dp}$ if $f=i$, and to $0.3\times\overline{\phi}_{i}^{dp}$ otherwise. For the NDFPP-Binomial and NDFPP-Normal, $\rho$ is defined as $1+0.3\times(|\mathcal{N_{F}}|-1)$. For the NDFPP-Normal, $\widetilde{\mu}_{fi}^{dp}$ is $\overline{\mu}_{f}^{dp}$ if $f=i$, and $0.3\times\overline{\mu}_{f}^{dp}$ otherwise. The parameter $\overline{\mu}_{f}^{dp}$ is given by $\nu_{\text{max}}\left[1-\frac{(P+1-p)(L+\hat{l}(f,d)+2)}{2(P+1)(L+1)}\right]$ for each protection level $p\in\mathcal{P}$, with $\hat{l}(f,d)$ returning zero for the $nd$ event, meaning that the mean increases with protection and decreases with intensity. The standard deviation $\widetilde{\sigma}^{d}_{f}$ is $\frac{1}{\rho}\left[\overline{\sigma}_{f}^{d}+0.3\sum_{i\in\mathcal{N}_{\mathcal{F}}\backslash\{f\}}\overline{\sigma}_{i}^{d}\right]$, where $\overline{\sigma}_{f}^{d}$ is $\nu_{\text{max}}\left[\frac{L+\hat{l}(f,d)+2}{8(P+1)(L+1)}\right]$. This standard deviation increases with intensity levels and is independent of protection, ensuring $4\overline{\sigma}_{f}^{d}\leq\overline{\mu}_{f}^{d1}\leq\cdots\leq\overline{\mu}_{f}^{dP}$.
Given that the costs $c_{f}^{p}$ are randomly generated, we consider five different seeds per instance configuration, (network, number of facilities, and capacity level). Thus, for NDFPP-Selection, NDFPP-Binomial, and NDFPP-Discrete, we have $5\times 2\times 3=30$ instances per seed (150 in total). For NDFPP-Normal, we have $5\times 2=10$ instances per seed (50 in total).
5.2.2 Comparative results for NDFPP-Selection.
For the case of discrete selection of distributions, the literature on endogenous uncertainty has mainly linearized the probability function (3) with different approaches. Here, we consider the DE-OP formulated in Bhuiyan et al. (2020) as an MILP. To solve this MILP, we implemented the same methodology as Bhuiyan et al. (2020): an L-shaped decomposition accelerated with nondominated cuts and valid inequalities. We compare this approach with the SAA-TP. Note that we cannot apply the SAA to the DE-OP as it has endogenous uncertainty.
[…]
0010101010202020203030303040404040505050506060606070707070808080809090909010010010010000101010102020202030303030404040405050505060606060707070708080808090909090100100100100Instances (%)SAA-TP $\left(\overline{gap}(\overline{x})(\%)\right)$DE-OP $\left(GAP(\%)\right)$
Comparison of the optimality gap obtained by the SAA-TP and the DE-OP.
Table 5.2.2 shows the experimental results of the 4- and 5-facility instances. In the table, $|K|$ and $|K_{t}|$ are the set of scenarios in the DE-OP (Program (2)) and in the deterministic equivalent of our transformed Program (4), respectively. Moreover, $GAP(\%)$ is the optimality gap for the DE-OP. For the DE-OP, all runs reached the time limit of two hours. We followed the procedure explained in Section 3.1 for the SAA-TP to obtain the feasible solution $\overline{x}$ and its estimators (see Appendix 7 for details). Even though the number of scenarios $|K_{t}|$ of the transformed program is extremely large (i.e.,$|\mathcal{D}|{|\mathcal{W}|}^{|\mathcal{N_{F}}||\mathcal{P}|}$), the number $|K|$ of scenarios of the DE-OP is small (i.e., $|\mathcal{D}|{|\mathcal{W}|}^{|\mathcal{N_{F}}|}$) for the considered instances. Thus, we can compute the real stochastic value of the feasible solution $\overline{x}$, i.e., $v(\overline{x})=c^{\top}\overline{x}+g(\overline{x})$. For the SAA-TP, we solve $M=50$ problems with $N=750$ sample scenarios because these values sufficed to obtain estimators with low variance and feasible, high-quality solutions. We present the over-estimator of the relative optimality gap, i.e., $\overline{gap}(\overline{x})(\%)={100\%\times[v(\overline{x})-\overline{v}_{M}^{N}]}/{v(\overline{x})}$. Because the lower-bound estimator $\overline{v}_{M}^{N}$ is a random variable, this over-estimator of the optimality gap may assume negative values. The computational time of the SAA-TP is in seconds. For all runs, the SAA-TP finished before the time limit of two hours. For comparison, we display the gap between the upper-bound values obtained by the SAA-TP and the DE-OP, i.e., $GAP_{UB}(\%)={100\%\times[UB-v(\overline{x})]}/{UB}$, and the gap between their lower bounds, i.e., $GAP_{LB}(\%)={100\%\times[LB-\overline{v}_{M}^{N}]}/{LB}$, where $UB$ and $LB$ are the upper and lower bounds for the DE-OP, respectively. Next, we summarize our main findings.
Result 1
The SAA-TP produces a feasible solution $\overline{x}$ with a considerably low over-estimator of the optimality gap, $\overline{gap}(\overline{x})(\%)$ is between $0\%$ and $1\%$ for all instances. The lower bound $\overline{v}_{M}^{N}$ and the optimality gap estimators have low variability, as shown by the standard deviation $\hat{\sigma}_{N,M}$ of the lower-bound estimator, which is inferior to $0.5\%$ of $\overline{v}_{M}^{N}$ for all instances (see Appendix 17 for details). For the considered instances, the computational time required by the SAA-TP was inferior to 2000s in 87% of the cases, with an average of around 990s.
Result 2
The SAA-TP’s over-estimator of the optimality gap is significantly smaller than the DE-OP’s optimality gap (the difference is around $24\%$ on average). The SAA-TP obtained these optimality gaps in significantly less time than the DE-OP, and, within the time limit, consistently found feasible solutions with slightly better values than those from the DE-OP.
Result 1 highlights the good performance of the SAA-TP. Result 2 shows that it outperformed the DE-OP. Figure 5.2.2, which shows the optimality gap for the DE-OP and the SAA-TP, corroborates this analysis. The SAA-TP obtained optimality gaps close to $0\%$ for all instances, whereas the DE-OP obtained optimality gaps inferior to $10\%$ only for $22\%$ of the instances. The DE-OP obtained optimality gaps of $10\%$$20\%$ for around $56\%$ of the instances, of $20\%$$30\%$ for around $8\%$ of the instances, and of $100\%$ for around $14\%$ of the instances. Even though the transformed Program (4) has more scenarios than the DE-OP, the SAA-TP solves different problem instances with fewer scenarios. Therefore, although solving the DE-OP is constrained to a relatively small number of scenarios, the transformed program can handle larger instances as we can use sampling techniques to address it. The SAA-TP obtained high-quality solutions by sampling only $N=750$ scenarios for each of the $M=50$ SAA problems, which, in most cases, is less than the number $|K|$ of scenarios of the DE-OP. Despite the number of scenarios, the SAA-TP solves mathematical programs smaller than the DE-OP (in terms of number of variables and constraints). The DE-OP’s MILP formulation contains numerous big-M constraints due to the linearization of the probability function (3), which leads to weak linear relaxations. In contrast, the transformation applied to obtain the NDFPP-Selection formulation (see Section 5.1.1) includes no extra variables or constraints (see Appendix 18 for a comparison of the number of variables and constraints in the DE-OP and the SAA-TP). Finally, although the SAA performed well in this case study, other sampling methods can be employed based on the problem’s specific requirements.
5.2.3 Analysis for NDFPP-Binomial, NDFPP-Discrete, and NDFPP-Normal.
In this section, we analyze the statistical estimators and feasible solution $\overline{x}$ of the SAA-TP for NDFPP-Binomial, NDFPP-Discrete, and NDFPP-Normal, which are MILPs (see Sections 5.1.2-5.1.4). In these cases, the DE-OP (Program (2) for discrete distributions) is nonlinear nonconvex, and the linearization of Bhuiyan et al. (2020) cannot be applied. Moreover, note that our transformed Program (4) for these variants has continuous random variables, making the explicit enumeration of scenarios impossible and requiring the implementation of sampling techniques. The DE-OP has a small number $|K|$ of scenarios for the NDFPP-Binomial and NDFPP-Discrete instances. Thus, we compute $v(\overline{x})$ for these cases. The number of scenarios of the DE-OP is infinite for NDFPP-Normal because the endogenous random variables are continuous. Consequently, we compute the estimator $\hat{v}_{N{{}^{\prime}}}(\overline{x})$ of $v(\overline{x})$. Tables 5.2.3 and 5.2.3 lists the experimental results of the 4- and 5-facility instances. The over-estimator of the optimality gap $\overline{gap}(\overline{x})(\%)$ is defined as in the previous section for NDFPP-Binomial and NDFPP-Discrete. For NDFPP-Normal, $\overline{gap}(\overline{x})(\%)={100\%\times[\hat{v}_{N{{}^{\prime}}}(\overline{x})-\overline{v}_{M}^{N}]}/{\hat{v}_{N{{}^{\prime}}}(\overline{x})}$. We set $M=50$ and $N=750$ for the same reason as in the previous section. For NDFPP-Normal, we use $N{{}^{\prime}}=5\times 10^{4}$. The SAA-TP finished for all instances before reaching 3 hours. The result is as follows:
Result 3
The SAA-TP obtained a feasible solution $\overline{x}$ with a significantly low statistical over-estimator of the optimality gap $\overline{gap}(\overline{x})(\%)$ between $-1\%$ and $1\%$ for all instances of the problem versions. The SAA’s estimators have high confidence, as shown by the standard deviation $\hat{\sigma}_{N,M}$ of the lower-bound estimator, which is inferior to $0.5\%$ of $\overline{v}_{M}^{N}$ for all cases, and the standard deviation $\hat{\sigma}_{N{{}^{\prime}}}(\overline{x})$ of the upper bound estimator, which is inferior to $0.2\%$ of $\hat{v}_{N{{}^{\prime}}}(\overline{x})$ for all NDFPP-Normal instances (see Appendix 19 for details). The computational time was, on average, around 1500s for NDFPP-Binomial, 3000s for NDFPP-Discrete, and 1200s for NDFPP-Normal.
[…]
Experimental results for NDFPP-Binomial and NDFPP-Discrete of the SAA-TP.
[…]
Experimental results for NDFPP-Normal of the SAA-TP.
[…]
The SAA-TP found high-quality solutions and estimators for the considered problems. The computational time of this approach for the NDFPP-Binomial, NDFPP-Discrete, and NDFPP-Normal exceeded that for the NDFPP-Selection. Solving the NDFPP-Discrete required the longest computational time. The transformation used in this case (i.e., inversion for discrete random variables; see Section 4.1.1) includes numerous binary variables and big-M constraints in the formulation. The initial formulation of NDFPP-Discrete contained bilinear terms, which were linearized by adding more big-M constraints in the formulation. For the same reason, the computational time of the SAA-TP for NDFPP-Binomial exceeded that for NDFPP-Selection. The computational time of the SAA-TP for solving NDFPP-Discrete increases with the number of capacity levels (possible realizations of the random variables). Therefore, a trade-off exists between how sample size affects the performance and confidence of the SAA’s statistical estimators and the associated computational time. For NDFPP-Normal, the number of evaluation scenarios $N{{}^{\prime}}$ is considerably large, which increases the computational time of the SAA-TP. Moreover, the standard deviation of the optimality gap is slightly larger than for the remaining versions because it includes the variability of $\hat{v}_{N{{}^{\prime}}}(\overline{x})$.
5.2.4 Value of the Stochastic Solution.
This section analyzes the relative VSS of all problem variants. The relative VSS is defined as ${(\textrm{{{{EEV}}}}-v^{*})}/{\textrm{{{{EEV}}}}}$, where $v^{*}$ is the optimal value of the recourse problem, and EEV is ${v(\hat{x}^{*})=c^{\top}\hat{x}^{*}+g(\hat{x}^{*})}$, with $\hat{x}^{*}$ the optimal solution of the EV. In Appendix 14, we detail the EVs (i.e., the deterministic problems with the average case scenario) according to the problem variants. We use the feasible solution $\overline{x}$ obtained by the SAA-TP to compute under estimators of the VSS, and we detail this computation in Appendix 20. We resume our main findings in Result 4.
Result 4
The SAA-TP obtained better solutions than the EV in all cases. Thus, solving the recourse problem instead of the EV is also advantageous in the context of endogenous uncertainty, where the distribution defined by $\hat{x}^{*}$ is not the same as that defined by $\overline{x}$ or $x^{*}$. Different instances of NDFPP-Selection showed a high VSS, around $30\%$ or more. Some instances of NDFPP-Binomial and NDFPP-Discrete also presented high VSSs, whereas others had VSSs of $1\%$$4\%$. Although the VSS for NDFPP-Normal approached $0\%$ for the 5-facility instances, it hovered around $3\%$$6\%$ for the 4-facility instances. Although the VSS is primarily instance-dependent, it is also affected by the problem (e.g., NDFPP-Selection and NDFPP-Normal for 15-node instances have considerably different VSSs). Finally, Appendix 20 contains the complete results.
6 Conclusions
We proposed a general method for addressing 2-SMIPD that uses random variable transformation. Our methodology (i) works for a general class of programs with endogenous uncertainty; (ii) obtains a stochastic program with exogenous uncertainty for which solution procedures are well-studied; (iii) yields MILPs or MICPs with exogenous uncertainty in many cases; (iv) works for both discrete and continuous endogenous random variables; and (v) is applicable to problems with independent and correlated random variables. We investigated techniques to define appropriate transformations of random variables to treat cases where the decision-dependent random variables follow classical probability distributions. We showed the existence of a simple transformation for the case of the discrete selection of distributions, which is the most commonly studied in previous research. We also demonstrated the proposed methodology by applying it to the NDFPP, a broadly studied problem. In the computational experiments, the SAA-TP produced feasible, high-quality solutions and high-confidence estimators in a short computational time.
Numerous research perspectives stem from this work. Future work could focus on (i) exploring additional techniques to define transformation functions; (ii) investigating approaches to deal with the current limitations of the method when $q(\omega)$, $W(\omega)$ and/or $H(\omega)$ are random, such as exploring valid inequalities for the discrete random variables inverse transformation; (iii) analyzing how endogenous uncertainty affects the definition of the VSS; (iv) examining stochastic programs with endogenous random variables for which the probability distribution is not known a priori; and (v) extending the NDFPP to account for assumptions investigated in previous research, such as post-disruption edge failure, protection (see, e.g., (Du and Peeta 2014)), and facility location (Li et al. 2021, 2022).
[…]
This research was funded by the SCALE-AI Chair in Data Driven Supply Chains, the FRQ-IVADO Research Chair in Data Science for Combinatorial Game Theory, and NSERC Grant No. 2024-04051. It was also enabled by support provided by Calcul Québec (www.calculquebec.ca) and the Digital Research Alliance of Canada (https://alliancecan.ca/).